### v15

# Investigation about a statement equivalent to Riemann Hypothesis (RH) applied to Dirichlet primitive L functions

ABSTRACT
We try to apply a known equivalence, for RH about Riemann $\zeta$ function, to Dirichlet L functions with primitive characters. The aim is to give a small contribution to the proof of the generalized version of Riemann Hypothesis (RH) ( i.e. GRH).
MSC-Class : 11M06, 11M26, 11M99
Keywords : Riemann Hypothesis ; Generalized Riemann Hypothesis ; Dirichlet L function ; Riemann Z functions ; Euler Product.
1. Introduction
Using Dirichlet characters $\chi(n)$, L functions, [3, p. 249, and 262], and, specially, ([6, p. 3, and 37] (whose conventions, i.e. $\chi_{0}$ as principal, is used throughout) are defined as :
(1.1) $\L (s.\chi)=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}=\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad,\quad\quad\quad\Re(s)>1$
The infinite product is called Euler Product, while $s:\ 0<\Re[s]<1$, and s:[s]=1/2s:\ $\Re$[s]=1/2 are called respectively: critical strip, and, critical line. The complex variable is $s=1/2+\epsilon+it$ throughout. GRH statement is:
“all the zeros inside critical strip of the expressions (1.1), or of their analytic continuations, are on critical-line”.
For reasons will be clear below we focus only on $\chi_{primitive}$, defined in [3, p. 168 -170]. Companion function like $\xi(s)$, ( [6, p. 62] ) for $\zeta(s)$, is $\xi(s,\chi_{primitive})$, ( [6, p. 71] ) for $L(s,\chi_{primitive})$, below their definition :
(1.2) $\xi(s)=\Gamma\left(\frac{s+2}{2}\right)(s-1)\frac{\zeta(s)}{\pi^{s/2}}\ \ \ \ ;\ \ \ \ \xi(s,\chi)=\left(\frac{q}{\pi}\right)^{\frac{s+\alpha}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)L(s,\chi)$
where $\Gamma(z)$ is defined in [3, p. 251], $q=congruence\quad modulus$. While $\alpha=0$ if $\chi_{primitive}(-1)=1$, even character, and, $\alpha=1$ if $\chi_{primitive}(-1)=-1$, odd character. Both (1.2 ) have the same zeros of the $\zeta(s)$ or $L(s,\chi)$ functions respectively in the critical strip [6] [7, p. 16]. Following [6, p. 71, and 79] the functional equation for $\xi(s)$ and $\xi(s,\chi)$ are:
(1.3) $\quad\xi(1-s)=\xi(s)\ \ \ ;\ \ \ \xi(1-s,\bar{\chi})=\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\xi(s,\chi)\quad with\quad\chi_{primitive}\ \ \quad$
Where $\bar{\chi}$ is the complex conjugate of $\chi$ and $\tau(\chi)$ is the gauss sum $\tau(\chi):=\sum_{m=1}^{q}\chi(m)e^{2\pi im/q}$ see [6, p. 65] , or [3, p. 165].
While zeros of $\zeta(s)$ have also a symmetry with respect to real axis, the $L(s,\chi)$ functions maintain this symmetry only for real characters $\chi$.
There is a link between, $\zeta(s)$, and, principal characters of congruence modulus $q$, see [3, p. 232]. It is reported below for easy reading:
(1.4) $L(s,\chi_{principal})=\zeta(s)\prod_{p\mid q}(1-p^{-s})$
So RH for $\zeta(s)$ (unproven until now) would prove also that Dirichlet functions $L(s,\chi)$ with principal characters have same zeros of $\zeta(s)$ inside critical strip. There is also a link between, $\psi$, primitive characters, and, character $\chi$, neither principal nor primitive ( see [3, p. 262] ), “inducing $\psi$ :
(1.5) $L(s,\chi)=L(s,\psi)\prod_{p\mid q}(1-\psi(p)p^{-s})\quad where\ \ \chi(n)=\chi_{principal}(n)\psi(n)$
To prove completely GRH for all characters of Dirichlet L functions (1.1), it is enough to prove RH for $\zeta(s)$,(question unsolved till now),and, to prove RH for $L(s,\chi_{primitive})$ . Here we focus only on second point. In particular only on odd primitive characters.
The structure of the paper is:
  • In section 2 we introduce functions useful for study phase behavior of $\xi(s,\chi_{primitive})$
  • In section 3 we show that these functions can be expressed by Euler Product , and, by P.N.T. we prove Theorem 1: $L(s,\chi_{odd\ primitive})$ fulfills RH at least for $|t|>T_{Asymp}(\alpha)$, where $T_{Asymp}(\alpha)$ is defined in (A.14 ).
  • in section 4 we find that the value of $T_{Asymp}(\alpha)$, for primitive characters, is surprisingly about $\pi$. We extend Theorem 1 for $|t|\leq T_{Asymp}(\alpha)$ where it applies.
  • In appendix A we give an unified treatment of derivatives along $t$ and $\epsilon$ in expressions where appears $\angle[\Gamma(z(s))]$ function excluding the $t=0$ neighborhood. The symbol $\angle[a]$ means phase of complex number $a$.
  • In appendix B we compare an useful approximation to (3.13 ).
  • In appendix C is justified the use of $L(s,\chi_{primitive})$ in critical strip.
  • In appendix D are discussed some observation from readers.
2. Angular Momenta and related Lemmas
The derivative of the phase of $\xi(s)$ for constant $\epsilon$ with respect to $t$ is :
$\ \ \ \ \ \frac{\partial}{\partial t}\arctan\left(\frac{\Im\xi(s)}{\Re\xi(s)}\right)$ $=$ $\frac{1}{1+\left(\frac{\Im\xi(s)}{\Re[\xi(s)]}\right)^{2}}\frac{\frac{\partial\Im\xi(s)}{\partial t}\Re\xi(s)-\frac{\partial\Re\xi(s)}{\partial t}\Im\xi(s)}{\left(\Re\xi(s)\right)^{2}}$
$=$ $\frac{\frac{\partial\Im\xi(s)}{\partial t}\Re\xi(s)-\frac{\partial\Re\xi(s)}{\partial t}\Im\xi(s)}{\left(\Re\xi(s)\right)^{2}+\left(\Im\xi(s)\right)^{2}}=\frac{\partial\angle[\xi(s)]}{\partial t}$
The numerator, that determines the sign of (2), can be seen as the angular momentum with respect to the origin of an unitary mass positioned in $(\Re\xi(s),\Im\xi(s))$ at time $t$ for constant $\epsilon$.
$DEFINITION\ 1:\ \ Angular\ Momentum\ for\ \xi(s,\chi)$
We can write also for $\xi(s,\chi)$:
(2.2) $\mathcal{L}[\xi(s,\chi)]:=\det\left(\begin{matrix}\Re[\xi(s,\chi)]&\Im[\xi(s,\chi)]\\ \frac{\partial}{\partial t}\Re[\xi(s,\chi)]&\frac{\partial}{\partial t}\Im[\xi(s,\chi)]\end{matrix}\right)=|\xi(s,\chi))|^{2}\times\frac{\partial\angle[\xi(s,\chi)]}{\partial t}\ \ :\ \ \chi_{primitive}$
2.1. LEMMA 1
If $A(s)$ is a derivable complex function and $F$ is a complex constant, then:
(2.3) $\mathcal{L}[FA(s)]=|F|^{2}\mathcal{L}[A(s)]$
PROOF: $\angle[F\ A(s)]=\angle[F]+\angle[A(s)]$ so $\frac{\partial\angle[F\ A(s)]}{\partial t}=\frac{\partial\angle[A(s)]}{\partial t}$. But $|FA(s)|^{2}=|F|^{2}|A(s)|^{2}$. From (2.2 ) follows (2.3).
[…]
Let us apply Lemma 1 to 1.3 ( remember $\chi_{primitive}$). For [6, p. 66] $|\tau(\chi)|^{2}=q$, then $\left|\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\right|=1$. But, at $\epsilon=0$ we have: $1-s=\bar{s}$. Besides, as $\Gamma(\bar{s})=\int_{0}^{\infty}x^{\bar{s}-1}e^{-x}dx=\bar{\Gamma}(s)$ ([7, p. 8]), $\left(\frac{q}{\pi}\right)^{\frac{\bar{s}+\alpha}{2}}=\bar{\left(\frac{q}{\pi}\right)^{\frac{s+\alpha}{2}}}$, and $L(\bar{s},\bar{\chi})=\bar{L}(s,\chi)$ ( (1.1)) , then for 1.2 ), and ( 1.3):
(2.4) $\xi(\bar{s},\bar{\chi})e^{-\frac{i}{2}\angle\left[\frac{i^{a}q^{\frac{1}{2}}}{\tau(\chi)}\right]}=\xi(s,\chi)e^{\frac{i}{2}\angle\left[\frac{i^{a}q^{\frac{1}{2}}}{\tau(\chi)}\right]}=\eta(t,\chi_{primitive})\ \ \in\Re\ \ ,\ \ \ \epsilon=0$
because if $\bar{z}=z$ then $z\in\Re$. So, posing $\eta^{\prime}(t,\chi)=\frac{d\eta(t,\chi)}{dt}$:
$\mathcal{L}[\xi(t,\epsilon=0,\chi_{primitive})]=\left|e^{-\frac{i}{2}\angle\left[\frac{i^{a}q^{\frac{1}{2}}}{\tau(\chi)}\right]}\right|^{2}\mathcal{L}[\eta(t)]=\left|e^{-\frac{i}{2}\angle\left[\frac{i^{a}q^{\frac{1}{2}}}{\tau(\chi)}\right]}\right|^{2}\det\left(\begin{matrix}\eta(t,\chi)&0\\ \eta^{\prime}(t,\chi)&0\end{matrix}\right)=0$
DEFINITION 2DEFINITION\ 2
Let us define ( see (1.3) ):
(2.5) $\eta(\frac{1}{2}+\epsilon+it,\chi)=\eta(s,\chi)=\left(e^{\frac{i}{2}\angle\left[\frac{i^{\alpha}k^{\frac{1}{2}}}{\tau(\chi)}\right]}\xi(s,\chi)\right);\ for\ \epsilon=0\quad it\quad is\quad real\quad\forall\chi_{primitive}$
Notice that for Lemma 1 $\forall s:\mathcal{L}[\xi(s,\chi_{primitive})]=\mathcal{L}[\eta(s,\chi_{primitive})]$ as $\left|e^{-\frac{i}{2}\angle\left[\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\right]}\right|^{2}=1\ \ \forall q$ and $\forall\chi_{primitive}$.
2.3. LEMMA 3
We have:
(2.6) $\left[\frac{\partial\mathcal{L}[\xi(s,\chi)])}{\partial\epsilon}\right]_{\epsilon=0}=\left[\frac{\partial\mathcal{L}[\eta(s,\chi)])}{\partial\epsilon}\right]_{\epsilon=0}=\eta(t,\chi)\left[-\frac{d^{2}\eta(t,\chi)}{dt^{2}}\right]+\left[\frac{d\eta(t,\chi)}{dt}\right]^{2}\ \ \;\ \ \ \chi=\chi_{primitive}$
PROOF:
$\left[\frac{\partial\mathcal{L}[\eta(s,\chi)])}{\partial\epsilon}\right]=\frac{\partial}{\partial\epsilon}det\left(\begin{matrix}\Re[\eta(\frac{1}{2}+\epsilon+it)]&\Im[\eta(\frac{1}{2}+\epsilon+it)]\\ \frac{\partial\Re[\eta(\frac{1}{2}+\epsilon+it)]}{\partial t}&\frac{\partial\Im[\eta(\frac{1}{2}+\epsilon+it)]}{\partial t}\end{matrix}\right)=$
(2.7) $\frac{\partial}{\partial\epsilon}\left[\Re[\eta(\frac{1}{2}+\epsilon+it)]\ \times\ \frac{\partial\Im[\eta(\frac{1}{2}+\epsilon+it)]}{\partial t}-\frac{\partial\Re[\eta(\frac{1}{2}+\epsilon+it)]}{\partial t}\ \times\ \Im[\eta(\frac{1}{2}+\epsilon+it)]\right]$
But for Cauchy-Riemann equations [5, p. 19], and, Lemma 2:
(2.8) $\frac{\partial\Im[\eta(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\eta(\frac{1}{2}+\epsilon+it)]}{\partial t}\ \ ;\ \ \Im[\eta(\frac{1}{2}+it)]=0\ \ ;\ \ \frac{\partial\Im[\eta(\frac{1}{2}+it)]}{\partial t}=0=\frac{\partial\Re[\eta(\frac{1}{2}+it)]}{\partial\epsilon}$
so from (2.8 ) we can equate , (for $\epsilon=0$), (2.7) to:
$\Re[\eta(\frac{1}{2}+it)]\ \times\ \frac{\partial^{2}\Im[\eta(\frac{1}{2}+it)]}{\partial\epsilon\partial t}-\frac{\partial\Re[\eta(\frac{1}{2}+it)]}{\partial t}\ \times\ \frac{\partial\Im[\eta(\frac{1}{2}+it)]}{\partial\epsilon}=$
$\eta(t,\chi)\left[-\frac{d^{2}\eta(t,\chi)}{dt^{2}}\right]+\left[\frac{d\eta(t,\chi)}{dt}\right]^{2}=[\eta^{\prime}(t,\chi)]^{2}-\eta^{\prime\prime}(t,\chi)\eta(t,\chi)$
So Lemma 3 is proved.
3. Lemmas on $L(s,\chi_{primitive})$ phase variations along $t$ computed by Euler product
If $\chi$ is not principal, as a primitive character, then $L(s,\chi)$ is an entire function ([3, p. 255] ), and, (1.1) converges absolutely for $\Re(s)>1$, while converges conditionally (see [1, p. 406] ) for $\Re(s)>0$ ( [10, p. 7], and, also unifomily [5, p. 7]). So we take $\Re[s]>0$ throughout.
In appendix C it is shown a proof for that.
The idea to exploit Euler product also inside critical strip is not new. See [9].
Here we wonder if , for primitive characters, it makes sense to think also of Euler product in $0<\Re(s)\leq 1$. We are interested in phase variations $\Delta\angle[L(s,\chi_{primitive})]$ computed through Euler product. We do not say that (1.1) is valid also in $0<\Re(s)\leq 1$, but that phase variations of Euler product makes sense for $\Re(s)>0$ in particular situations treated below.
3.1. Euler product in critical strip
Let us consider the expressions:
$\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad\ \ \;\ \ \ \ L(s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}$
It is known that $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad\$can be seen as the product of infinite geometrical serie, each one, with common ratio $\frac{\chi(p_{j})}{p_{j}^{s}}$, ( with $\left|\frac{\chi(p_{j})}{p_{j}^{s}}\right|<1$ ) i.e. $\frac{1}{1-\frac{\chi(p)}{p^{s}}}=\sum_{n=0}^{\infty}\left(\frac{\chi(p_{j})}{p_{j}^{s}}\right)^{n}$. Let us consider only a finite number of primes $j_{max}:\ p_{1},p_{2}...p_{max}\ ,\ i.e.j_{max}=\pi(p_{max})$, and so also a finite numbers of geometrical series to be multiplied. Besides notice that geometrical series can be written by a finite number of terms:
We can multiply $j=1,...j_{max}=\pi(p_{max})$ geometric series with infinite terms, but we can alternatively choose to consider, in each geometrical series, the sum beyond a certain exponent $\alpha_{j}=\left\lceil\frac{\ln(p_{max})}{\ln(p_{j})}\right\rceil$ as a whole. i.e.
$\frac{1}{1-\frac{\chi(p_{j})}{p_{j}^{s}}}=\sum_{n=0}^{\infty}\left(\frac{\chi(p_{j})}{p_{j}^{s}}\right)^{n}=\sum_{n=0}^{\alpha_{j}-1}\left(\frac{\chi(p_{j})}{p_{j}^{s}}\right)^{n}+\left(\frac{\chi(p_{j})}{p_{j}^{s}}\right)^{\alpha_{j}}(1-\chi(p_{j})p_{j}^{-s})^{-1}$
So we can have a finite number of terms for each of the $j=1...j_{max}$ geometric series instead of infinite terms.
For example as an $n\approx p_{max}$ cannot have two divisors both $\sqrt{p_{max}}$ then all the series with $p_{j}$ from $\approx\sqrt{p_{max}}$ to $p_{max}$ are simplified as :
$1+\frac{\chi(p_{j})}{p_{j}^{s}}+\left(\frac{\chi(p_{j})}{p_{j}^{s}}\right)^{2}\frac{1}{1-\frac{\chi(p_{j})}{p_{j}^{s}}}\ \ ;\ \sqrt{p_{max}}<p_{j}\leq p_{max}$
At the end we get:
$EP(t,\epsilon,\chi,p_{max})=$
(3.1) $\prod_{p=2}^{p_{max}}\frac{1}{1-\frac{\chi(p)}{p^{s}}}=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}+\sum_{\rho>p_{max}}\frac{\chi(\rho)}{\rho^{s}}\prod_{\alpha_{j}>0}(1-\chi(p_{j})p_{j}^{-s})^{-1}=L(s,\chi,p_{max})+R(s,\chi,p_{max})$
$Where\ \rho=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}...p_{max}^{\alpha_{max}}\ ;\alpha_{j}=0,1,2...\ \ and\ \ j_{max}=\pi(p_{max})$
In (3.1) we can have huge amount of terms :
(3.2) $N^{o}(terms)=3\prod_{j=1}^{j_{max}-1}\left(\left\lceil\frac{\ln(p_{max})}{\ln(p_{j})}\right\rceil+1\right)\ \ ;\ \ last\ factor\left\lceil\frac{\ln(p_{max})}{\ln(p_{max})}\right\rceil+1\ \ ;\ \ is\ put\ to\ 3$
For example for $p_{max}=31$ we have in (3.2 ) that $N^{o}(terms)=787320$, whoose only 31 belong to $L(s,\chi,p_{max})$. The others belong to $R(s,\chi,p_{max})=R(t,\epsilon,\chi,p_{max})$.
So we does not need to care, in applying Euler rearrangement for $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, about absolute convergence of the $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime$, because we have an huge, but, finite amount of terms.
We have:
(3.3) $R(t,\epsilon,\chi,p_{max})=\sum_{\rho>p_{max}}\frac{\chi(\rho)}{\rho^{s}}\prod_{\alpha_{j}>0}(1-\chi(p_{j})p_{j}^{-s})^{-1}=\ \ ;\ \ Where\ \rho=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}...p_{{j_{max}}}^{\alpha_{j_{max}}}>p_{max}\ ;\ \alpha_{j}=0,1,2$...
$\sum_{\rho>p_{max}}[\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}\prod_{\alpha_{j}>0}(1-\chi(p_{j})p_{j}^{-s})^{-1}$
Where $j_{max}=\pi(p_{max})$ and the:
$N^{o}$ of addends of $R(t,\epsilon,\chi,p_{max})$ is : $3\prod_{j=1}^{j_{max}-1}\left(\left\lceil\frac{\ln(p_{max})}{\ln(p_{j})}\right\rceil+1\right)\ \ -p_{max}$.
Notice that variation in $\Delta t$ is due almost completely to $\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]$, because with $p_{j}<<\rho$ we have $\ln(p_{j})<<\ln(\rho)$, and for big $p_{j}\leq p_{max}$, for $\epsilon>0$, we have $|\chi(p_{j})p_{j}^{-s})^{-1}|<<1$ for $p_{j}$ big enough. Besides we are interested in $\Delta t$ small.
Only first term in right part of (3.1), i.e $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, tend to (C.1), so, due to (3.3), we have, considering the whole critical strip:
(3.4) $\L (s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}\neq EP(t,\epsilon,p_{max}\rightarrow\infty)=\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad,\quad\quad\quad 0<\Re(s)<1\ \ \$
This is immediate by considering that left part can be zero in critical strip while each factor of right part cannot. Notice however that our aim is to use $\angle[L(s,\chi)]$, and, also the phase is undefined when $L(s,\chi)=0$.
But if we consider the difference (or better the phase difference) in a finite interval, suitably far from $\L (s.\chi)=0$, the conclusion is different. Let us take:
[…]
If $\Delta t=t_{2}-t_{1}$ is suitable small, i.e. $p^{*}$ suitable big, from (3.1) we have:
$\Delta[L(s.\chi)]=L(t_{2},\epsilon.\chi)-L(t_{1},\epsilon.\chi)=\Delta[L(s.\chi,p_{max})]\ ;\ p_{max}\rightarrow\infty$
Let us consider the identity:
(3.6) $\Delta\left[R(t,\epsilon,\chi,p_{max})\right]_{\Delta t}=\int_{t_{1}}^{t_{2}}\frac{\partial R(t,\epsilon,\chi,p_{max})}{\partial t}dt$
where in (3.3) we have an huge but finite number of terms with zero mean value. If $p_{max}>>p^{*}$ we have a lot of cancellation in $\Delta t$ interval for single (3.6) addend. The cancellation will be complete, for a particular $\rho$, only if
(3.7) $\Delta t\times\ln(\rho)=\frac{2\pi}{\ln(p^{*})}\ln(\rho)=m\times 2\pi\ ;\ m\in N\ ,\ i.e.\ \rho=(p^{*})^{m}$
When condition (3.7) is not verified the derivative of a term of (3.1, in (3.6) integral, reaches same phase in $t_{1}$ not in $t_{2}$, but in a closer point $t_{2}+\Delta t_{\rho}$ where $\Delta t_{\rho}$ is positive or negative, and $\Delta t_{\rho}=0$ if (3.7) hold.
Such $\Delta t_{\rho}$ become smaller in absolute value, and, also the amplitude of oscillation evidenced in (3.3), decreases as $p_{max}$ ( and, cosequently $\rho$ ) grows.
So the error terms at $\Delta t$ extrema must go to zero as $p_{max}\rightarrow\infty$, even if extremely slowly, because all $\ln(\rho)\times t$ are different.
In order to improve convergence , pointing to bigger addends of (3.3), (i.e. at $\rho$ close to $p_{max}$ ), we can choose:
(3.8) $p_{max}=(p^{*})^{m};\ m\in N\ ,\ m\rightarrow\infty$
Let us consider $(R(t,\epsilon,\chi,p_{max})$ terms with $\rho=p_{1}^{\alpha_{1}}...p_{max}^{\alpha_{max}}=p_{max}\times r$, and $\frac{\ln(r)}{\ln(p_{max})}$ small. These terms give an almost null contribution to the differences from $t_{2}$ to $t_{1}$ of (3.1 ), i.e. $L(s,\chi,p_{max})+R(s,\chi,p_{max})$ for $n$ and $\rho$ close to $p_{max}$ because phase difference in $\Delta t$ is: $\ln(\rho)\Delta t=\ln(p_{max})\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)\ \frac{2\pi}{\ln(p_{max})}=2\pi\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)$ , and for $p_{max}\rightarrow\infty$ it tends to $2\pi$. So as concern these small $r$ terms we have (considering mean values in $\Delta t$ interval):
(3.9) $\left.\Delta\angle[EP(s.\chi,p_{max})]\right|_{\Delta t}\rightarrow\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}\ \ ;$
where $p^{*}$ is whatever big.
So condition (3.8) is an hint to convergence at lower $p_{max}$ while the argument stemming from identity (3.6) can be applied to generic $p_{max}>>p^{*}$ condition, but, it requires higher $p_{max}$ values to reach a comparable convergence. In a simplistic view we could say:
(3.3) is mainly an oscillation with period close to $\frac{2\pi}{\ln(p_{max})}$, if we take $\Delta t=m\times\frac{2\pi}{\ln(p_{max})}\ ;\ m\in N$ we filter away the main part of (3.3) .
Notice that the mean phase variation in $\Delta t$ is :
(3.10) $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}$
3.2. Application of Euler Product in critical strip
Inspired by (3.9) we can consider, for $s:\ L(s,\chi)\neq 0$, that:
$\angle\left[\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\right]=-\sum_{\forall p}\angle\left[1-\chi(p)p^{-s}\right]=-\sum_{\forall p}\angle\left[1-\chi(p)p^{-1/2-\epsilon-it}\right]=$
$\sum_{\forall p}\angle\left[\{p^{1/2+\epsilon}-\chi(p)p^{-it}\}p^{-1/2-\epsilon}\right]=$
$\sum_{\forall p}\angle\left[p^{+1/2+\epsilon}-\{\cos(-\ln(p)t+\angle[\chi(p)])+i\sin(-\ln(p)t+\angle[\chi(p)])\}\right]=$
$\sum_{\forall p}\angle\left[\left(p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])\right)+i\left(\sin(\ln(p)t-\angle[\chi(p)])\right)\right]$
So the phase computed from Euler product in (1.1) till $p_{max}\rightarrow\infty$ is:
(3.11) $\sum_{p<p_{max}\ :\ gcd(p,q)=1}\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\rightarrow\angle[EP(s,\chi_{primitive},p_{max})]\ \ \ ;\ \ \Re(s)>0$
where $q$ is the congruence modulus.
Far from $(t,\epsilon):\ L(t,\epsilon,\chi)=0$ we can apply (3.9).
Expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$, where total phase is involved, converges for $\Re(s)>0$, but, in singular points.They are, see (2 ) , $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}$,where $\xi(s^{*}_{k},\chi_{primitive})=L(s^{*}_{k},\chi_{primitive})=0$ (see (1.2 ), and (A.2) ). We neglect, for the moment, double zeros. To be specific $\xi(s,\chi)$ is equal ( [6, p. 82]) to the infinite Hadamard product, that has single factors like
(3.12) $1-\frac{s}{s^{*}_{k}}=1-\frac{(0.5+\epsilon^{*}_{k})^{2}+tt^{*}_{k}+i(0.5+\epsilon^{*}_{k})(t-t^{*}_{k})}{(0.5+\epsilon^{*}_{k})^{2}+(t^{*}_{k})^{2}}=(t^{*}_{k}-t)\gamma\ ;\ \gamma\in C$
So, close to the single zero $(t_{k}^{*},\epsilon_{k}^{*})\$, along $\epsilon$ with $0.5<\epsilon\leq\epsilon^{*}_{k}\geq 0$ line, $\angle[\xi(s,\chi)]$ has, at $t=t^{*}_{k}$, an abrupt phase jump of $\pi$, due to the factor (3.12). And in (1.2 ), this abrupt phase change, can be attributed only to $\angle[L(s,\chi)]$. See appendix A. But singular points are also on half-line $0.5<\epsilon<\epsilon^{*}_{k}\ ,\ t=t_{k}^{*}$. Here there is a total phase discontinuity of, at least, $2\pi$. Because, as $\delta\rightarrow 0$, from $(\epsilon<\epsilon_{k}^{*}\ ,\ t=t^{*}_{k}-\delta)$ to $(\epsilon<\epsilon^{*}\,\ t=t^{*}_{k}+\delta)$ there is a total phase jump of, at least, $2\pi$, i.e., the phase change in circling one or two zeros couterclockwise ( argument principle [5, p. 25] ). Notice this discontinuity is not present in (2 ), where total phase is not used.
Now let us see how to evaluate phase variations of expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$ using (3.9 ).
In order to trade resolution along $t$ with computational burden it is convenient to address to mean phase variation in an interval $(t_{1},t_{2})$ . Besides in computing the mean phase difference in the interval $\Delta t=\frac{2\pi}{\ln(p^{*}}$ (3.5) for $p_{max}\rightarrow\infty$ the oscillations of $R(t,\epsilon,p_{max})$ (3.3) are rejected.
We consider, $t$ increment in order to compute incremental ratio with respect to $t$, and, define the value $p^{*}$ like in (3.5).
For $\epsilon>-\frac{1}{2}$ (i.e. for $\Re(s)>0$) we can operate this way: we choose $p^{*}$ big at will, afterward we compute (3.13) till $p_{max}\rightarrow\infty$, or, better $m\rightarrow\infty$. We can iterate these operations with a greater $p^{*}$. In shortcut $\{p_{max}=(p^{*})^{m}\rightarrow\infty\}p^{*}\rightarrow\infty]$, better $\{p_{max}=(p^{*})^{m}\ ,\ m\rightarrow\infty\}p^{*}\rightarrow\infty]$. Notice that we do not interchange differentiation with infinite sum. Instead we compute the mean incremental ratio of the “almost everywhere” differentiable function $\angle[L(s.\chi_{primitive})]$ using (3.9) between $t_{1}$ and $t_{2}$; i.e. $\frac{\Delta\angle[L_{EP}(s,\chi_{primitive})]}{\Delta t}$ where $\Delta t=\frac{2\pi}{\ln(p^{*}}$ using all primes. Afterward every iteration to let $p^{*}\rightarrow\infty$, requires new computation of mean incremental ratio.
(3.13) $\left[-\frac{\ln(p^{*})}{2\pi}\sum_{p=2}^{p=p_{max}=(p^{*})^{m}\ ,\ m\rightarrow\infty\ :\ gcd(p,q)=1}\left\{\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\right]_{p^{*}\rightarrow\infty}=$
$\left\{\left[\frac{\Delta\angle[L_{EP}(s,\chi_{primitive})]}{\Delta t}\right]_{p_{max}=(p^{*})^{m}\ ,\ m\rightarrow\infty}\right\}_{p^{*}\rightarrow\infty}\rightarrow\ \ \ \frac{\partial\angle[L(s,\chi_{primitive})]}{\partial t}\ \ ;\ \ \{\Re(s)>0\}\setminus\{sing.points\}$
It happens that , for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, (3.13) grows without bounds not only on the zeros, but also on the segment $0.5<\epsilon<\epsilon^{*}_{k}\ ,\ t=t_{k}^{*}$, where a step discontinuity arises in total phase of (3.11). So, when computing (3.13), we find a kind of ridge stemming from the $k^{th}$ zero $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it_{k}^{*}\ \ ,\ 0.5>\epsilon^{*}_{k}\geq 0$ toward negative reals. Notice that for $p$ big ( i.e. $p^{1/2+\epsilon}>>1\geq\left|\cos(\ln(p)t-\angle[\chi(p)])\right|\ \ \$), the content of braces summed in (3.13 tends to be:
(3.14) $\left\{\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\rightarrow$
$\left\{\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\rightarrow\left\{\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}=$
$\frac{\sin\left(\ln(p)\left\{t+\frac{\pi}{\ln(p^{*})}\right\}-\angle[\chi(p)]\right)-\sin\left(\ln(p)\left\{t-\frac{\pi}{\ln(p^{*})}\right\}-\angle[\chi(p)]\right)}{p^{1/2+\epsilon}}=$
(3.15) $\frac{2\ \cos(\ln(p)t-\angle[\chi(p)])\sin\left(\frac{\pi\ln(p)}{\ln(p^{*})}\right)}{p^{1/2+\epsilon}}$
When there is a zero in $s=s^{*}=1/2+\epsilon^{*}+it^{*}$, then (3.13) grows without bounds for $s\rightarrow(s^{*})^{+}$ , with constant $t=t^{*}$. It must grow without bounds, for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, also for $0.5<\epsilon<\epsilon^{*}$ with $t=t^{*}$.
$PROOF.$
We have that (3.13) grows without bound for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, for $s\rightarrow(s^{*})^{+}$ also if we start the sum (3.13) from $P_{1}>>2$ .
Moreover for $0.5<\epsilon<\epsilon^{*}\ ,\ t=t^{*}$, as $p^{1/2+\epsilon}-\cos[\ln(p)t-\angle[\chi(p)]>0\ ,\ \forall p\geq 2$, $\forall t$ and $\forall\epsilon>-1/2$, (3.14) written as (3.15) shows that $\forall p>P_{1}$, where $P_{1}$ is a prime big enough, the content in braces in (3.13) is bigger, because multiplied by $p^{\epsilon^{*}-\epsilon}$, and preserving sign (as $\epsilon>-1/2$).
We can use the below reported Abel summation formula [1, p. 407] for $p\geq P_{1}$. Let $\{a_{k}\}$ and $\{b_{k}\}$ be two sequences of complex numbers, and let $A_{n}=\sum_{k=1}^{n}a_{k}$, then $\sum_{k=1}^{n}a_{k}b_{k}=A_{n}b_{n+1}+\sum_{k=1}^{n}A_{k}(b_{k}-b_{k+1})$. We want to start from a big prime $P_{1}$ so we pose:
$A_{n}=\sum_{k=1}^{n}a_{k}=\sum_{p_{k}=P_{1}}^{p_{n}}\frac{2\ \cos(\ln(p_{k})t-\angle[\chi(p)])\sin\left(\frac{\pi\ln(p_{k})}{\ln(p^{*})}\right)}{p_{k}^{1/2+\epsilon^{*}}}\rightarrow\infty$ for $p_{n}\rightarrow\infty\ ,\ \forall P_{1}$. Besides $b_{k}=p_{k}^{\epsilon^{*}-\epsilon}$, with $p_{1}=P_{1}$. As $b_{n+1}=-[(0-b_{1})+(b_{1}-b_{2})+..+(b_{n}-b_{n+1})]$, we can write: $\sum_{k=1}^{n}a_{k}b_{k}=A_{n}b_{n+1}+\sum_{k=1}^{n}A_{k}(b_{k}-b_{k+1})=$ $A_{n}b_{1}+\sum_{k=1}^{n}[A_{n}(b_{k+1}-b_{k})-A_{k}(b_{k+1}-b_{k})]=$ $A_{n}b_{1}+\sum_{k=1}^{n}[(A_{n}-A_{k})(b_{k+1}-b_{k})]$ $\left[\sum_{k=1}^{n}a_{k}\right]b_{1}+\sum_{k=1}^{n}[\{\sum_{k_{1}=k+1}^{n}a_{k_{1}}\}(b_{k+1}-b_{k})]$. For hypothesis on $\{a_{k}\}$ and $\{0<b_{k}<b_{k+1}\}$ we have : $\left[\sum_{k=1}^{n}a_{k}\right]b_{1}\rightarrow\infty$ for $n\rightarrow\infty$. Also $\{\sum_{k_{1}=k+1}^{n}a_{k_{1}}\}(b_{k+1}-b_{k})]\rightarrow\infty$ for $n\rightarrow\infty\ ,\ \forall k$. So (3.13) must grow without bounds also for $0.5<\epsilon<\epsilon^{*}$ with $t=t^{*}$, and, $p_{n}\rightarrow\infty$.
ENDofPROOFEND\ of\ PROOF
So we find, again, the ridge for (3.13) from $s_{k}^{*}$, non-trivial zeros of $\L (s.\chi_{primitive})$, toward negative reals, but, by a different approach. The PNT in arithmetic progression, with $q$ as congruence modulus [3, p. 148], states that for each $h$-class of primes, ($p\equiv h\pmod{q}$, $gcd(h,q)=1$, and $0<h<q$) :
(3.16) $\pi(x)_{h,q}\sim\frac{Li(x)}{\phi(q)}\quad;\quad\forall h:gcd(h,q)=1\ ,and\ \ Li(x)=\int_{2}^{x}\frac{dt}{\ln(t)}$
Where $\pi(x)_{h,q}$ is the count of $h$-class primes till $x$, and, aritmetic function $\phi(q)$ [3, p. 25], is the Euler totient function. There are $\phi(q)$ classes of primes $p\equiv h\pmod{q}$ where $gcd(h,q)=1$, and, $0<h<q$. For all primes in a particular $h$-class, $\angle[\chi(p)_{p\equiv h\pmod{q}\ ,\ gcd(h,q)=1\ ,\ 0<h<q}]=\angle[\chi(h)]$ is constant. For non principal characters (in particular for primitives ones) ( see [3, p. 256]) we have:
[…]
We want to point out at $h$-prime classes in (3.13). If we limit to a finite $p_{max}$ we can write (3.13 ) as the sum of the $\phi(q)$ sums involving the $h$-prime classes:
(3.18) $\frac{\ln(p^{*})}{2\pi}\sum_{h<q\ :\ gcd(h,q)=1}\left\{\sum_{p<p_{max}\ :\ p\equiv h(mod\ q)}\left[\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right]_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\right\}$
In order to take $p_{max}=(p^{*})^{m}\rightarrow\infty$ we have to fill all the sums in natural order so that partial sums are not affected. Grouping with braces must not alter the order of the sum because there is not absolute convergence for $0<\Re(s)<1$. In other words, it is true that rearranging terms we can change the limit of a conditional converging series at will, [1, p. 411], but, if we rearrange within a finite number of terms (i.e. till $p_{max}$) generated by natural order, then, the usual rules of addition must hold, like in (3.1). Notice that in (3.18), in the sum in braces $\angle[\chi(p)]=\angle[\chi(h)]$ because it deals only with one $h$-prime class.
3.3. LEMMA 4: use of logarithmic integral
The sum of the $\phi(q)$ integrals $h$-dependent like (3.19) is null $\forall p_{max}$:
(3.19) $\frac{\ln(p^{*})}{\pi}\sum_{h<q:gcd(h,q)=1}\frac{1}{\phi(q)}\int_{2}^{p_{max}}\frac{\cos(\ln(y)\ t-\angle[\chi(h)])}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]=0\quad\forall p_{max}$
PROOF. Each $h:0<h<q$, defines an $h$-class of primes modqmod\ q
$\sum_{h<q:gcd(h,q)=1}\frac{1}{\phi(q)}\int_{2}^{p_{max}}\frac{\cos(\ln(y)\ t-\angle[\chi(h)])}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]=$
$\sum_{h<q:gcd(h,q)=1}\cos(\angle[\chi(h)])\frac{1}{\phi(q)}\int_{2}^{p_{max}}\frac{\cos(\ln(y)\ t)}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]+$
$\sum_{h<q:gcd(h,q)=1}\sin(\angle[\chi(h)])\frac{1}{\phi(q)}\int_{2}^{p_{max}}\frac{\sin(\ln(y)\ t)}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]=$
Integrals are not $h$-dependent and so can be put into common factor obtaining:
$\frac{\int_{2}^{p_{max}}\frac{\cos(\ln(y)\ t)}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]}{\phi(q)}\sum_{h<q:gcd(h,q)=1}\cos(\angle[\chi(h)])+$
$\frac{\int_{2}^{p_{max}}\frac{\sin(\ln(y)\ t)}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(y)}{\ln(p^{*})}\right)d[Li(y)]}{\phi(q)}\sum_{h<q:gcd(h,q)=1}\sin(\angle[\chi(h)])$
But for (3.17)
$\sum_{h<q:gcd(h,q)=1}\sin(\angle[\chi(h)])=0\ \ \ \ \ \ \ \ ,and,\ \ \ \ \ \sum_{h<q:gcd(h,q)=1}\cos(\angle[\chi(h)])=0$
So $\forall p_{max}$ (3.19) is null independently of the values assumed by integrals at common factor. Notice , always complying with the natural order of the sum in (3.18) till $p_{max}$, we could distribute (only for indexes less than running index) the $\phi(q)$ integrals in (3.19) into the $\phi(q)$ braces of (3.18). If $\theta(z,h,t)=\ln(z)t-\angle[\chi(h)]$, we get by (3.5):
(3.20) $\frac{\ln(p^{*})}{2\pi}\sum_{h<q\ :\ gcd(h,q)=1}$
$\left\{\sum_{p<p_{max}\ :\ p\equiv h(mod\ q)}\left[\arctan\left(\frac{\sin(\theta(p,h,t))}{p^{1/2+\epsilon}-\cos(\theta(p,h,t))}\right)\right]_{t_{1}}^{t_{2}}-\frac{\int_{2}^{p_{max}}\frac{2\cos(\theta(x,h,t))}{y^{1/2+\epsilon}}\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)d[Li(x)}{\phi(q)}\right\}$
as usual $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$, and, the overall result does not change, provided all operations for $p<p_{max}$ and $x<p_{max}$ are carried out at same time, while no computations with greater $p_{max}$ is just done. This is true because (3.13) converges $\forall s\setminus\{singular\ points\}$, and, for Lemma 4 the contribution of (3.19) till whatever $p_{max}$ is always zero. Then is applicable [1, p. 385 Theorem 10.2]: if two series of complex terms converges, then converges also their sum to the sum of their limits.
3.4. LEMMA 5: Euler Product on critical line
We want to compute $\left[\frac{\partial\angle[\xi(s,\chi)]}{\partial t}\right]$ around $\epsilon=0$ for $t\neq t^{*}_{k}$.
Neglecting minor contributions, from appendix A, in particular (A.2), (A.14), and (A.15), we have:
$\left[\frac{\partial\angle[\xi(s,\chi)]}{\partial t}\right]_{\epsilon\approx 0\ ,\ t\neq t^{*}_{k}}=$
(3.21) $\left[\frac{\partial\angle[L(s,\chi_{primitive})]}{\partial t}\right]_{\epsilon=0}+\epsilon\times\left[\frac{\partial^{2}\angle[L(t,\epsilon=0,\chi_{primitive})]}{\partial\epsilon\partial t}\right]_{\epsilon=0}+\ln\left[\sqrt{\frac{tq}{2\pi}}\right]\quad\pm\frac{\epsilon}{4t^{2}}\ \ ;\ \ \forall t\neq t^{*}_{k}$
Where $\pm$ refers rispectively to odd and even characters (A.15). We want to use (3.13) to compute previous equation for: $\epsilon=0$. In order to avoid singular points we choose for $\epsilon\leq\epsilon^{*}_{k}$ $\quad|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}\$, while this condition can be dropped for $\epsilon>\epsilon^{*}_{k}$. So we can write ($t>T_{Asymp}(\alpha)$ (A.14) ):
(3.22) $\ln\left[\sqrt{\frac{t\times q}{2\pi}}\right]+\left\{\left[\frac{\Delta\angle[L_{EP}(s,\chi_{primitive})]}{\Delta t}\right]_{p_{max}\rightarrow\infty}\right\}_{p^{*}\rightarrow\infty}\rightarrow\left[\frac{\partial\angle[\xi(t,\epsilon,\chi)]}{\partial t}\right]_{\epsilon=0}=0$
PROOF. From Lemma 2, appendix A and (3.13 ) follows (3.22 ). End of proof.
In half-plane $\Re[s]>0$ (see also figg. 5, 4, and , 3)) , we can remark that:
  • (1) $\mathcal{L}[\xi(s,\chi)]$ is defined $\forall s$, see (2.2),
  • (2) $\frac{\partial\angle[\xi(t,\epsilon,\chi)]}{\partial t}$ is defined everywhere but points $s^{*}_{k}=1/2+\epsilon^{*}_{k}+it^{*}_{k}$ where $L(s^{*}_{k},\chi)=0$. See (2), where, only phase difference between vector of real and imaginary components with respect to vector of their $t$-derivative matters.
  • (3) $\frac{\Delta\angle[L_{EP}(t,\epsilon,\chi)]}{\Delta t}$ instead is defined everywhere but points $(\epsilon\leq\epsilon^{*}_{k}\ ,t=t^{*}_{k})$ , see (3.13 ). Here is used total phase.
3.5. LEMMA 6: Sign estimate of $\frac{\partial^{2}\angle[L(s,\chi_{primitive})]}{\partial\epsilon\partial t}$
Suppose $\epsilon>0$ but arbitrarily small, $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}$, with (3.5), and $p^{*}$ big (fixed), and $p_{max}\rightarrow\infty$. Then,for (3.22) we have that (3.13) evaluated in $\epsilon=0$ is tending to $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]<0$ as $(p_{max}\rightarrow\infty)p^{*}\rightarrow\infty]$’, in order to verify Lemma 2. We affirm that the same (3.13), evaluated for small $\epsilon>0$ is always negative but with lower absolute value, so that (3.22 ) has a positive increment for small $\epsilon>0$. i,e (at least for odd primitive characters, see (3.21) ):
(3.23) $\left|\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right|>\left|\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}\right|$
with $\left\{\left[\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi_{primitive})]}{\Delta t}\right]_{p_{max}\rightarrow\infty}\right\}_{p^{*}\rightarrow\infty}\rightarrow-\ln\left[\sqrt{\frac{t\times q}{2\pi}}\right]$
PROOF . Our aim is difference between (3.13) and (3.19), i.e. (3.20), computed in natural order. For Lemma 4 the result for $\{p_{max}\rightarrow\infty\}p^{*}\rightarrow\infty]$ does not change from (3.13) that converges apart singular points. See (3.11).
Let us analyze the peculiar oscillations in (3.20) integrands and, to simplify, suppose, at first, that $x<p^{*}$, i.e $\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)>0$. Considering that the zero crossing of $\cos[\ln(x)t-\angle[\chi(h)]]$ are at $\ln(x)t-\angle[\chi(h)]=2k\pi\mp\frac{\pi}{2}\rightarrow\ln(x)=\frac{2k\pi\mp\frac{\pi}{2}+\angle[\chi(h)]}{t}$, let us see zero transitions (subscript $0t$ means ‘zero transition’).
  • Zero transition at increasing $\cos[\ln(x)t-\angle[\chi(h)]]$ :
    (3.24) $\quad x_{0t}(k,h)=e^{(2\pi k-\pi/2+\angle[\chi(h)])/t}\quad increasing\quad cosine$
  • Zero transition at decreasing $\cos[\ln(x)t-\angle[\chi(h)]]$ :
    (3.25) $\quad x^{\prime}_{0t}(k,h)=e^{(2\pi k+\pi/2+\angle[\chi(h)])/t}\quad decreasing\quad cosine$
Corresponding points to $x_{0t}(k+1,h),x^{\prime}_{0t}(k,h)$ defined on (3.19) can be defined on (3.13) : $y_{0t}(k+1,h),y^{\prime}_{0t}(k,h)$, thinking (3.14) as continuous function of $y=p$. In more detail: let us choose $\phi(q)$ functions of $y$, continuous variable, in place of discrete $p$:
(3.26) $F(y,\chi(h),\epsilon)=\left\{\arctan\left(\frac{\sin(\ln(y)t-\angle[\chi(h)])}{p^{1/2+\epsilon}-\cos(\ln(y)t-\angle[\chi(h)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}$
$y_{0t}(k+1,h),y^{\prime}_{0t}(k,h)$ are defined as positive slope zero transitions and negative slope zero transition respectively (similarly to $x_{0t}(k+1,h),x^{\prime}_{0t}(k,h)$)
Besides for $k$ big (i.e. $p$ or $y$ big) $x_{0t}(k+1,h)\rightarrow y_{0t}(k+1,h)$ and $x^{\prime}_{0t}(k,h)\rightarrow y^{\prime}_{0t}(k,h)$ as we want. To build (3.13) minus (3.19), i.e (3.20), then, negative contributions for (3.19) become positive contributions for final result.
Negative contribution for (3.19) ( i.e. positive contribution for final result) is given by the (3.19) integral in the intervals $\Delta x_{k,h}$ :
(3.27) $x^{\prime}_{0t}(k,h)-x_{0t}(k,h)=x_{0t}(k,h)[e^{(\pi/t)}-1]=\left(\frac{\pi}{t}+\frac{\pi^{2}}{2t^{2}}+..\right)e^{(2\pi k-\pi/2+\angle[\chi(h)])/t}=\Delta x_{k,h}\approx\frac{\pi}{t}\hat{x}_{0t}(k,h)$
where $\hat{x}_{0t}(k,h)$ is a suitable value inside $\Delta x_{k,h}$
(3.28) $O^{+}_{k,h}(\epsilon)_{\int dLi}=\left|\frac{\ln(p^{*})}{\pi\phi(q)}\int_{x>x_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{x<x^{\prime}_{0t}(k,h)}\frac{\cos[\ln(x)t-\angle[\chi(h)]]\sin[\pi\ln(x)/\ln(p^{*})]}{\sqrt{x}x^{\epsilon}}dLi(x)\right|$
Negative contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}$):
(3.29) $O^{-}_{k,h}(\epsilon)_{\sum}=\left|\frac{\ln(p^{*})}{2\pi}\sum_{p>y_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{p<y^{\prime}_{0t}(k,h)}F(p,\chi(p),\epsilon)\right|$
Positive contribution for (3.19) (i.e. negative contribution for final result) is given by the (3.19) integral in the intervals $\Delta x_{k,h}^{\prime}$ (see (3.27) ) :
(3.30) $x_{0t}(k+1,h)-x^{\prime}_{0t}(k,h)=x^{\prime}_{0t}(k,h)[e^{(\pi/t)}-1]=\left(\frac{\pi}{t}+\frac{\pi^{2}}{2t^{2}}+..\right)e^{(2\pi k+\pi/2+\angle[\chi(h)])/t}=\Delta x_{k,h}^{\prime}\approx\frac{\pi}{t}\hat{x}^{\prime}_{0t}(k,h)$
(3.31) $O^{-}_{k,h}(\epsilon)_{\int dLi}=\left|\frac{\ln(p^{*})}{\pi\phi(q)}\int_{x>x^{\prime}_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{x<x_{0t}(k+1,h)}\frac{\cos[\ln(x)t-\angle[\chi(p)]]\sin[\pi\ln(x)/\ln(p^{*})]}{\sqrt{x}x^{\epsilon}}dLi(x)\right|$
Positive contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}^{\prime}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}^{\prime}$)
(3.32) $O^{+}_{k,h}(\epsilon)_{\sum}=\left|\frac{\ln(p^{*})}{2\pi}\sum_{p>y^{\prime}_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{p<y_{0t}(k+1,h)}F(p,\chi(p),\epsilon)\right|$
Until now we have dealt with $p_{max}<p^{*}$. If we let $p^{*}$ fixed and $p_{max}$ increases without bound we have that in the intervals :
(3.33) $(p^{*})^{m}<p_{max}<(p^{*})^{m+1}$
if $m$ is even, the factor $\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)$ in (3.19) is positive, so, the contributions in corresponding $x$-intervals are as above. In other words, the sign contributions of final result, i.e. (3.13 ) minus (3.19), or , (3.20), in intervals $\Delta x,\Delta x^{\prime},\Delta y,\Delta y^{\prime}$, is as that stated in ( 3.29), (3.32) , and, (3.28), (3.31). Instead where $m$, in (3.33), is odd, then, in every interval above the sign changes concurrently. There are also intermediate situations, but, we will see they can be neglected. So we could compute (3.13) minus (3.19), till $p_{max}$, ( interchanging $O^{+}\leftrightarrow O^{-}$ when $m$ is odd in (3.33) ) as in :
(3.34) $\sum_{k}\sum_{h}\left[O^{+}_{k,h}(\epsilon)_{\int dLi}+O^{+}_{k,h}(\epsilon)_{\sum}\right]-\sum_{k}\sum_{h}\left[O^{-}_{k,h}(\epsilon)_{\sum}+O^{-}_{k,h}(\epsilon)_{\int dLi}\right]=\sum_{+}^{p_{max}}(\epsilon)-\sum_{-}^{p_{max}}(\epsilon)$
calling all negative contributions till $p_{max}$:
(3.35) $\sum_{-}^{p_{max}}(\epsilon)=\sum_{k}\sum_{h}O^{-}_{k,h}(\epsilon)_{\sum}+O^{-}_{k,h}(\epsilon)_{\int dLi}$
and all positive contributions till $p_{max}$:
(3.36) $\sum_{+}^{p_{max}}(\epsilon)=\sum_{k}\sum_{h}O^{+}_{k,h}(\epsilon)_{\sum}+O^{+}_{k,h}(\epsilon)_{\int dLi}$
We stress again, (3.34 ) is computed without changing the order given by $p$ or $x=y$, like in (3.18 ) , i.e. partial sums are unaffected. Limit difference in (3.34 ) is in form: $\infty$-$\infty$, and, we know that if $|t-t_{k}^{*}|>\frac{2\pi}{\ln(p^{*})}$ limit result is finite . In other words: positive quantities: $O^{+}_{kh}(\epsilon)_{\int dLi}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\int dLi}$,and , $O^{+}_{k,h}(\epsilon)_{\sum}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\sum}$ like in ( 3.29), and (3.32) , and, (3.28), (3.31) are only “place-holders” filled according to progress of $x=y<p_{max}$, and with chosen sign following the $m$ parity in (3.33). Consider now the ratios :
(3.37) $\frac{O^{-}_{k,h}(\epsilon^{\prime})_{\sum}}{O^{+}_{k,h}(\epsilon^{\prime})_{\int dLi}}\ \ \ \ \ \ ;\ \ \ \ \ \ \frac{O^{+}_{k,h}(\epsilon^{\prime})_{\sum}}{O^{-}_{k,h}(\epsilon^{\prime})_{\int dLi}}$
for each $h$ and for $k$ big, defined respectively in intervals $\Delta y_{k,h}\rightarrow\Delta x_{k,h}$ ( see (3.28) and (3.29 ) ) , and in intervals $\Delta y_{k,h}^{\prime}\rightarrow\Delta x_{k,h}^{\prime}$ ( see (3.31) and (3.32) ) .
LEMMA 7: $\forall h,k$ as $k\rightarrow\infty$ both (3.37 ) tend to unity.
PROOF. PNT ratio in small intervals [8, p. 23] assures that:
(3.38) $\pi(x+\Phi(x))-\pi(x)\sim\frac{\Phi(x)}{\ln(x)}\ \ with\ \ x\rightarrow\infty\ \ ,if\ \Phi(x)\geq x^{7/12-\epsilon(x)}$
where $\epsilon(x)$ in (3.38) tend to zero as $x\rightarrow\infty$. Imposing this on equi-sign intervals, for (3.30) and (3.27):
(3.39) $\Delta x_{k,h}=\frac{\pi}{t}x>x^{7/12}\rightarrow\frac{\pi}{t}x^{5/12}>1\rightarrow x>\left(\frac{t}{\pi}\right)^{12/5}$
i.e. for $x$ big enough the distribution of primes in equal-sign intevals $\Delta x_{k,h}\rightarrow\Delta y_{k,h}$, and, for (3.16) also $h$-class primes, can be close to theorethical one as we want. In [17, p. 65] it is dealt about the sum of prime sampled functions: “in the neighborhood of the number $x$ the average density of the primes is $1/\ln(x)$. On this basis, if one should wish an estimate for the sum of $F(p)$ over all primes $p<x$, the natural approximation would be”
(3.40) $\sum_{p<x}F(p)\sim\int_{2}^{x}\frac{F(t)dt}{\ln(t)}=\int_{2}^{x}F(t)d[Li(t)]\rightarrow\frac{\sum_{p<x}F(p)}{\int_{2}^{x}F(t)d[Li(t)]}\rightarrow 1\ \ for\ \ x\rightarrow\infty$
PNT theorem is same statement of (3.40) with $F(p)=1$.
We use then (3.40) in intervals (3.30) and (3.27) for $k\rightarrow\infty$, with $F(p)$ function as in (3.15) or as in (3.26). For $p,x,y,k\rightarrow\infty$ (3.15) and (3.26) become the same function in intervals $\Delta x\rightarrow\Delta y\ \ ,\ \ \Delta x^{\prime}\rightarrow\Delta y^{\prime}$ that grow without bouds, fulfillig so (3.39). So thesis is proved. Besides if $\Delta x_{k,h}\approx\Delta y_{k,h}$, and, (3.39) is fulfilled, the $h$-class primes are $\approx\frac{\Delta x}{\phi(q)\ln(\hat{x})}$, so, far from $p_{max}\approx mp^{*}$, neglecting $|\cos()|$ and $|\sin()|$ factors from (3.15), or equivalent shaping from (3.26), and, with $\hat{x}$ as a suitable value inside $\Delta x_{k,h}$, $O^{\pm}_{k,h}$ can be evaluated as:
(3.41) $O^{\pm}_{k,h}\approx\frac{\pi}{t}\frac{2\hat{x}^{0.5}}{\phi(q)\ln(\hat{x})}$
[…]
(3.42) $\frac{\sum_{+}^{p_{max}}(\epsilon^{\prime\prime})}{\sum_{-}^{p_{max}}(\epsilon^{\prime})}=\frac{\sum_{k}\sum_{h}O^{+}_{k,h}(\epsilon^{\prime\prime})_{\int dLi}+O^{+}_{k,h}(\epsilon^{\prime\prime})_{\sum}}{\sum_{k}\sum_{h}O^{-}_{k,h}(\epsilon^{\prime})_{\sum}+O^{-}_{k,h}(\epsilon^{\prime})_{\int dLi}}=\rho(\epsilon^{\prime\prime}=\epsilon^{\prime}=0,p_{max}\rightarrow\infty)\ \ \ \rightarrow\ \ 1\ \ \,$
This is true in a continuous way also considering that $p_{max}=[p^{*}]^{m}$ and $m\rightarrow\infty$. Besides equal-sign intervals at higher $p$, or, $k$, have greater weight then equal-sign intervals at lower $p\approx x,$ or $y$ . See (3.41). So the rare zero transitions of factor $\sin\left(\frac{\pi\ln(x)}{\ln(p^{*})}\right)$ in $\Delta x_{h,k}$ are without influence in the ratio (3.42). The differences between $\Delta x$ and $\Delta y$ fades away with big $x=y\approx p$, so they do not influence the ratio (3.42). Besides when $p_{max}$ satisfy (3.33) with odd $m$ then in ratio (3.42) we add $O^{-}_{k,h}(\epsilon^{\prime\prime})_{\sum}+O^{-}_{k,h}(\epsilon^{\prime\prime})_{\int dLi}$ at numerator and $O^{+}_{k,h}(\epsilon^{\prime})_{\int dLi}+O^{+}_{k,h}(\epsilon^{\prime})_{\sum}$ at denominator, and, we have always the ratio beween all positive versus all negative contributions.
Now we want the ratio of all positive contributions to final result (3.34 ) at different $\epsilon$ value. In detail : $\forall h<q\ ,gcd(h,q)=1$, $\epsilon^{\prime\prime}>0$, and $\epsilon^{\prime}=0$:
(3.43) $\frac{\sum_{+}^{p_{max}}(\epsilon)}{\sum_{+}^{p_{max}}(0)}=\frac{\sum_{k}\sum_{h}O^{+}_{k,h}(\epsilon)_{\int dLi}+O^{+}_{k,h}(\epsilon)_{\sum}}{\sum_{k}\sum_{h}O^{+}_{k,h}(0)_{\int dLi}+O^{+}_{k,h}(0)_{\sum}}=\rho^{+}(\epsilon,p_{max})\ <\\ 1;\ \ \epsilon>0$
and the same for negative contributions ( we do not mention the interchange we do when $m$ is odd in (3.33)). Besides in intermediate cases; i.e. when $p_{max}\approx(p^{*})^{m}$, we could have a split in intervals $\Delta x,\Delta y$ or $\Delta x^{\prime},\Delta y^{\prime}$, that, in any case, do not alter the limit of (3.42) or (3.43) ):
(3.44) $\frac{\sum_{-}^{p_{max}}(\epsilon)}{\sum_{-}^{p_{max}}(0)}=\frac{\sum_{k}\sum_{h}O^{-}_{k,h}(\epsilon)_{\sum}+O^{-}_{k,h}(\epsilon)_{\int dLi}}{\sum_{k}\sum_{h}O^{-}_{k,h}(0)_{\sum}+O^{-}_{k,h}(0)_{\int dLi}}=\rho^{-}(\epsilon,p_{max})\ \ \ <\ \ 1;\ \ \epsilon>0$
(3.44 ) and (3.43 ), if $\epsilon>0$ is fixed and $p_{max}\rightarrow\infty$ goes to zero, because if $\epsilon>0$ and $k$ i.e. $p_{max}\rightarrow\infty$, then $p^{-\epsilon}\rightarrow 0$. But we have $\forall p_{max}$:
(3.45) $\sum_{+}^{p_{max}<\infty}(\epsilon>0)<\sum_{+}^{p_{max}<\infty}(\epsilon=0)\ \ \ \ \ ;\ \ \sum_{-}^{p_{max}<\infty}(\epsilon>0)<\sum_{-}^{p_{max}<\infty}(\epsilon=0)$
Besides the indeterminate limit (3.34 ), of $\infty$-$\infty$ form, for $\epsilon>0$ ($p^{*}$ is big and fixed), gives :
(3.46) $\sum_{-}^{p_{max}\rightarrow\infty}(\epsilon)=\sum_{+}^{p_{max}\rightarrow\infty}(\epsilon)+X(\epsilon,p_{max})\ \ ;\ \ X(\epsilon=0,p_{max}\rightarrow\infty)=-\left[\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi_{primitive})]}{\Delta t}\right]_{p_{max}\rightarrow\infty}$
Last expression, for $p^{*}\rightarrow\infty\ ,t\neq t^{*}_{k}$, approaches the smooth function $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$.
For big $p_{max}<\infty$, $\exists\epsilon>0$ suitable small (for $\epsilon=0$ the ratios $\rightarrow 1$), so that:
[…]
The (3.47) ratios approach $1$ from below for $\epsilon\rightarrow 0^{+}$, and, are close as we want, i.e.:
(3.48) $\rho^{+}(\epsilon,p_{max})=\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}\rightarrow\rho^{-}(\epsilon,p_{max})=\frac{\sum_{-}^{p_{max}<\infty}(\epsilon>0)}{\sum_{-}^{p_{max}<\infty}(\epsilon=0)}$
So, as $p_{max}$ grows, $\rho^{+}(\epsilon,p_{max})$ and $\rho^{-}(\epsilon,p_{max})$ can be close to each other as we want, as (3.42) holds also for $\epsilon^{\prime}=\epsilon^{\prime\prime}>0$. In other words we choose an $\epsilon^{\prime}$ small enough and a $p_{max}^{\prime}<\infty$, then we compute the $\rho^{\pm}(\epsilon^{\prime},p_{max}^{\prime})$. If they are too low with respect to $1$ we choose anoter $\epsilon^{\prime\prime}<\epsilon^{\prime}$ and we try again if this is satisfying for us . . and so on. We stress that there is no interchange of limits between $\epsilon\rightarrow 0$ and $p_{max}\rightarrow\infty$. We can write the identity:
(3.49) $X(\epsilon>0,p_{max}<\infty)=\left\{\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}\right\}\left[\sum_{+}^{p_{max}<\infty}(\epsilon=0)\right]-$
$\left\{\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)+X(\epsilon>0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)+X(\epsilon=0,p_{max})}\right\}\left[\sum_{+}^{p_{max}<\infty}(\epsilon=0)+X(\epsilon=0,p_{max})\right]$
Referring to (3.46), (3.47), and, (3.48), for big $p_{max}$,
$X(\epsilon>0,p_{max})=\left\{\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)+X(\epsilon>0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)+X(\epsilon=0,p_{max})}\right\}X(\epsilon=0,p_{max})=\rho^{-}(\epsilon,p_{max})X(\epsilon=0,p_{max})$ matches with:
$\left[\sum_{+}^{p_{max}<\infty}(\epsilon=0)\right]\left\{\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}-\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)+X(\epsilon>0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)+X(\epsilon=0,p_{max})}\right\}=$
$\left[\sum_{+}^{p_{max}<\infty}(\epsilon=0)\right]\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}\left\{1-\frac{1+\frac{X(\epsilon>0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}}{1+\frac{X(\epsilon=0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}}\right\}=$
(3.50) $\left[\sum_{+}^{p_{max}<\infty}(\epsilon=0)\right]\frac{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}\ \ \ \frac{\frac{X(\epsilon=0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}-\frac{X(\epsilon>0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon>0)}}{1+\frac{X(\epsilon=0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}}=$
$\rho^{+}(\epsilon,p_{max})\left(X(\epsilon=0,p_{max})-\frac{X(\epsilon>0,p_{max})}{\rho^{+}(\epsilon,p_{max})}\right)\frac{1}{1+\frac{X(\epsilon=0,p_{max})}{\sum_{+}^{p_{max}<\infty}(\epsilon=0)}}=0$
As $p_{max}$ grows, $\rho^{+}(\epsilon,p_{max})\rightarrow\rho^{-}(\epsilon,p_{max})$ for suitable small $\epsilon>0$, both approach 1, but, from below. See (3.47) and ( 3.48). So $X(\epsilon>0,p_{max})=\rho^{\pm}(\epsilon,p_{max})X(\epsilon=0,p_{max})$ for $p_{max}$ big enough. Then we can affirm that (3.50) is zero, and, identity (3.49) forces $X(\epsilon>0,p_{max})=-\left.\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ to reach from below $\left.\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right|_{p^{*}\ fixed,\ p_{max}}$ as $\epsilon\rightarrow 0^{+}$ for big $p_{max}$. This is (3.23), that holds $\forall p^{*}$ big. So we can state :
Corollary of Lemma 6. From (3.46 ) and from (3.13) we can write for $\epsilon\rightarrow 0^{+}$:
(3.51) $\frac{1}{\epsilon}\left[\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}-\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right]_{\ p_{max}\rightarrow\infty,\ p^{*}\ big}\rightarrow\left[\frac{\partial^{2}\angle[L(s,\chi_{primitive}))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\geq 0$
PROOF. The difference in square brackets is between mean values in $t\pm\frac{\Delta t}{2}$ (3.5) of same function with different $\epsilon$, that in $|t-t^{*}_{k}|>\frac{2\pi}{\ln(p^{*})}\ ,\ \epsilon\rightarrow 0$ is smooth in $t$, (i.e. $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$ for $\epsilon=0$), so, for $p^{*}$ big the difference of the mean in the intervals (3.5) is close to the function on the right in (3.51). Besides (3.23) holds for every big $p^{*}$, so, we can work on (3.46 ). For [2, p. 186 (Cauchy condition for series)] : a series $\sum a_{n}=X(\epsilon,p_{max})$ converges iff $\forall\delta>0\ \exists P:\forall n>P$ we have: $|\sum_{n>P}^{n=p+P}a_{n}|<\delta\ \ \forall p\geq 1$ ( i.e. the difference to the limit is $\delta$).
If (3.51) is false, i.e. $\frac{(-X(\epsilon>0,p_{max}\rightarrow\infty))-(-X(\epsilon=0,p_{max}\rightarrow\infty))}{\epsilon}<0$, for $\epsilon\rightarrow 0^{+}$, then if: $|X(\epsilon>0,p_{max}\rightarrow\infty)-X(\epsilon=0,p_{max}\rightarrow\infty)|=5\delta$,
we can find a certain $\epsilon^{\prime}$ and $p^{\prime}_{max}$: $|X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon^{\prime}>0,p_{max}\rightarrow\infty)|<\delta$,and, $|X(\epsilon=0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p_{max}\rightarrow\infty)|<\delta$. So we must have : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)-X(\epsilon=0,p^{\prime}_{max}<\infty)|>3\delta$ , i.e : $X(\epsilon^{\prime}>0,p^{\prime}_{max}<\infty)>X(\epsilon=0,p^{\prime}_{max}<\infty)$, i.e. $\rho^{\pm}(\epsilon^{\prime},p^{\prime}_{max})>1$. But this is not possible for $p^{\prime}_{max}$ big, and $\forall p^{*}$, see (3.45 ), (3.47). See Lemma 6 (3.23). So (3.51) must be $\geq 0$.
3.6. LEMMA 8 : $\left[\frac{\partial\mathcal{L}[\xi(t,\epsilon,\chi_{odd\ primitive})])}{\partial\epsilon}\right]_{\epsilon=0}>0\ ,\forall t\neq t^{*}_{k}\ ,\forall|t|>T_{Asymp}(\alpha)$
For $T_{Asymp}(\alpha)$ see (A.14 ). Referring to (2.5) we have:
(3.52) $\forall t\ \ with\ \ \eta(t,\chi)\neq 0\rightarrow\left(\frac{d\eta(t,\chi)}{dt}\right)^{2}-\eta(t,\chi)\times\frac{d^{2}\eta(t,\chi)}{dt^{2}}>0\ \ where\ \ \eta(t,\chi)=\eta(1/2+it,\chi),\ and\ \chi_{odd\ primitive}$
PROOF.
From definition (2.2), and, (2.5):
$\mathcal{L}[\eta(s,\chi)]=\mathcal{L}[\xi(s,\chi)]=|\xi(s,\chi))|^{2}\times\frac{\partial\angle[\xi(s,\chi))]}{\partial t}=|\eta(s,\chi))|^{2}\times\frac{\partial\angle[\eta(s,\chi))]}{\partial t}\ \ :\ \ \chi_{primitive}$
From (3.22 ) , as for Lemma 2:
$\left[\frac{\partial\angle[L(s,\chi_{primitive})]}{\partial t}\right]_{\epsilon=0}=-\ln\left(\sqrt{\frac{tq}{2\pi}}\right)\ ;\ \forall t\neq t^{*}_{k}$
we have that, from (1.2 ), (A.13), (A.15) and (3.21) :
$\left[\frac{\partial^{2}\angle[\xi(s,\chi)]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\ =\left[\frac{\partial^{2}\angle[L(s,\chi_{primitive})]}{\partial\epsilon\partial t}\right]_{\epsilon=0}+\left[\frac{\partial^{2}\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial\epsilon\partial t}\right]_{\epsilon=0}=$
(3.53) $\left[\frac{\partial^{2}\angle[L(s,\chi_{primitive})]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\quad\pm\frac{1}{4t^{2}}\ \ ;\ \ \forall t\neq t^{*}_{k}\ \ ;+\ odd\ \ ;-\ \ even\ \ \chi$
On the other hand:
$\left[\frac{\partial\mathcal{L}[\xi(s,\chi)]}{\partial\epsilon}\right]_{\epsilon=0}=\left\{\frac{\partial|\xi(s,\chi)|^{2}}{\partial\epsilon}\times\frac{\partial\angle[\xi(s,\chi))]}{\partial t}\right\}_{\epsilon=0}+\left[|\xi(s,\chi))|^{2}\times\frac{\partial^{2}\angle[\xi(s,\chi))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\ \ :\ \ \chi_{primitive}$
but, the expression in braces $\left\{\frac{\partial|\xi(s,\chi)|^{2}}{\partial\epsilon}\times\frac{\partial\angle[\xi(s,\chi))]}{\partial t}\right\}_{\epsilon=0}=0$ because for $t\neq t^{*}_{k}$: $\left\{\frac{\partial\angle[\xi(s,\chi))]}{\partial t}\right\}_{\epsilon=0}=0$, and $\left\{\frac{\partial|\xi(s,\chi)|^{2}}{\partial\epsilon}\right\}_{\epsilon=0}$ is finite ( as $\xi(s,\chi)$ is holomorphic). The same is true if we replace $\xi$ by $\eta$.
So, referring only to odd primitive characters ( see (A.15) ), and, using corollary to Lemmas 6 (3.51) that states $\left[\frac{\partial^{2}\angle[L(s,\chi_{odd\ primitive}))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\geq 0$, then, (3.53), implies that: $\left[\frac{\partial^{2}\angle[\eta(s,\chi))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}=\left[\frac{\partial^{2}\angle[\xi(s,\chi))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}>0$ (for $t>T_{Asymp}(\alpha)$, and $t\neq t^{*}_{k}$). So also (2.5):
$\left\{\frac{\partial}{\partial\epsilon}\mathcal{L}[\xi(s,\chi_{primitive})]\right\}_{\epsilon=0}=\left\{\frac{\partial}{\partial\epsilon}\mathcal{L}[\eta(s,\chi_{primitive})]\right\}_{\epsilon=0}=$
(3.54) $\left\{|\eta(s,\chi_{primitive})|^{2}\frac{\partial}{\partial\epsilon}\frac{\partial}{\partial t}\angle[\eta(s,\chi_{primitive})]\right\}_{\epsilon=0}>\delta>0\ \ ;\ \ \forall t:t\neq t^{*}_{k}$
Where $\delta=|\eta(t,\epsilon=0,\chi_{primitive})|^{2}\times\left(\frac{1}{4t^{2}}\right)>0$. END of proof.
LEMMA 9 : $\left[\frac{\partial\mathcal{L}[\xi(t,\epsilon,\chi_{odd\ primitive})])}{\partial\epsilon}\right]_{\epsilon=0}>0$ also $\forall t=t^{*}_{k},\ if\ \epsilon^{*}_{k}>0\ ,\forall|t|>T_{Asymp}(\alpha)$ .
[…]
3.7. THEOREM 1 : GRH is true for the odd primitive L-functions except for at most finitely many exceptions whose imaginary parts live in a finite interval $|t|<T_{Asymp}(\alpha)$
In [4, p. 6] is reported a statement as equivalent to R.H. for $\zeta(s)$: ”The Riemann hypothesis is equivalent to the statement that all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative.”. Let us see why. Suppose we are close to extremal points. Then $\Re[\xi(s)]$, is an harmonic function, obeying to Laplace equation $\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon^{2}}+\frac{\partial^{2}\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t^{2}}=0$, with null total curvature. See (2.8). As $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}$, and $\Im[\xi(\frac{1}{2}+it)]=0$, then extremal points on critical line (that are saddle points) are first candidate to look for off-critical line zeros. So, at extremal points (relative maxima end minima), if the curvature is toward $\epsilon,t]$ plane along $t$, it will be in the opposite direction along $\epsilon$. Simmetrically if the curvature is toward $\epsilon,t]$ plane along $\epsilon$, it will be in the opposite direction along $t$. Only in second case off-critical line zeros are possible because we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]<\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are tending toward the plane $\epsilon,t]$ along $\epsilon$. Instead in first case we have for example $\Re[\xi(t_{0},\epsilon_{0}=0)]>0$ and $\Re[\xi(t,\epsilon_{0}+\Delta\epsilon>0)]>\Re[\xi(t_{0},\epsilon_{0}=0)]$, i.e. we are increasing distance from plane $\epsilon,t]$ along $\epsilon$. In first case the curvature along $t$ must be negative, i.e. $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$. In second case we have instead $\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$ so $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}>0$. Notice the sign of the product $\xi(t,\epsilon=0)\frac{d^{2}\xi(t,\epsilon=0)}{dt^{2}}<0$ is invariant also if we choose $\Re[\xi(t_{0},\epsilon_{0}=0)]<0$ if all local maxima of $\xi(t,\epsilon=0)$ are positive and all local minima are negative. This means that if all extremal points of $\Re[\xi(t_{0},\epsilon_{0}=0)]$ comply with this rule, no off-critical line zero is possible because $\frac{\partial\Im[\xi(\frac{1}{2}+\epsilon+it)]}{\partial\epsilon}=-\frac{\partial\Re[\xi(\frac{1}{2}+\epsilon+it)]}{\partial t}\ \$. So $|\Im[\xi(\frac{1}{2}+\epsilon+it)]|$ grows with $\epsilon$ far from $\Re[\xi(\frac{1}{2}+it)]$ extremal points. We do not say that the set of points of $(t,\epsilon)$ plane where $\Im[\xi(s)]=0$ are perpendicular to critical line at $t_{0}$ where $\Re[\xi(\frac{1}{2}+it_{0})]$ is extremal, but, the truth is not far from it [12, p. 3] , and [13, p. 9,fig 6].
General facts of holomorphic behavior are present also for $\eta(t,\epsilon,\chi_{primitive})$, besides it is real on critical line as $\xi(t,\epsilon)$ so [4, p. 6] equivalence must be valid also for it. Besides for Lemma 8 and Lemma 9 we have: $\eta^{\prime}(t,\chi)]^{2}-\eta(t,\chi)\eta^{\prime\prime}(t,\chi)=\left\{\frac{\partial}{\partial\epsilon}\mathcal{L}[\xi(s,\chi)]\right\}_{\epsilon=0}>0\ ,\forall t:\ \eta(t,\chi)\neq 0$. But this means that at extremal points, i.e. $\frac{d\eta(t^{\prime},\chi)}{dt}=0$, local maxima are positive and local minima negative because:
(3.55) $\eta(t^{\prime},\chi)\left[\frac{d^{2}\eta(t^{\prime},\chi)}{dt^{2}}\right]<0$
So for $\eta(t,\epsilon=0,\chi)=0$ the zero ( or zeros if double) are on critical line. But also for $\eta(t,\epsilon=0,\chi)\neq 0$ expression (3.55) implies that relative maxima and minima must comply with no zeros off-critical line. So, for Lemma 8 and Lemma 9 , RH is granted at least in $|t|>T_{Asymp}(\alpha)$, see (A.14 ). So for a given $\L (s,\chi_{odd\ primitive})$ only a finite numbers of off-critical line zeros exist in $|t|<T_{Asymp}(\alpha)$, because the function is analytic in $|t|<T_{Asymp}(\alpha)$, if, there were infinitely many zeros there, it would be an accumulation point and the function would be identically zero, a contradiction.
4. Toward the proof $\forall t$ and $\forall$ modulus $q$ of primitive (odd) characters
With argument in appendix A and Lemmas of previous section, we proved RH for odd primitive characters in $|t|>T_{Asymp}(\alpha)$. In other words if $|t|>T_{Asymp}(\alpha)$ we are allowed to write (A.13) i.e. $\frac{\partial\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}=\ln\left(\sqrt{\frac{tq}{2\pi}}\right)+O(t^{-2})$, and $O(t^{-2})\rightarrow+(-1)^{\alpha+1}\frac{\epsilon}{4t^{2}}$ for $\epsilon\approx 0$ ($\alpha=0$ even, and $\alpha=1$ odd characters).
Here we find same results of appendix A without Stirling formula like in expression (A.5). The purpose is to evaluate (A.13) in the $t=0$ neighborhood. The starting point is the formula ([7, p. 8]) )
(4.1) $\Gamma(z)=\lim_{N\to\infty}\frac{N!\ \ (N+1)^{z}}{z(z+1)...(z+N)}$
where $z=\frac{s+\alpha}{2}=\frac{1/2+\epsilon+it+\alpha}{2}$ (see (1.2), and (A.4´) ) .
For $N\rightarrow\infty$ we sum and subtract: $\sum_{n=1}^{N}\frac{1}{n}\rightarrow\ln(N)+\gamma\approx\ln(N+1)+\gamma$, where $\gamma=0.577216..\ \$is the Euler - Mascheroni constant.
(4.2) $\Im[\ln(\Gamma(z))]=\angle[\Gamma(z)]=\angle[z\ln(N+1)]-\angle[z]-\angle[z+1]...-\angle[z+N]=$
$\left(\ln(N+1)-\sum_{1}^{N}\frac{1}{n}\right)\frac{t}{2}-\arctan\left(\frac{\frac{t}{2}}{\frac{1/2+\epsilon+\alpha}{2}}\right)+\sum_{1}^{N}\left\{\frac{t}{2n}-\arctan\left(\frac{\frac{t}{2n}}{1+\frac{1/2+\epsilon+\alpha}{2n}}\right)\right\}=$
$\gamma\frac{t}{2}-\arctan\left(\frac{\frac{t}{2}}{\frac{1/2+\epsilon+\alpha}{2}}\right)+\sum_{1}^{N}\left\{\frac{t}{2n}-\arctan\left(\frac{\frac{t}{2n}}{1+\frac{1/2+\epsilon+\alpha}{2n}}\right)\right\}$
Let us fix NN in (4.1 ), and, in (4.2), then if we take derivative with respect to $t$:
$\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}=\frac{-\gamma}{2}-\left(\frac{1}{1/2+\epsilon+\alpha}\right)\left(\frac{1}{1+\left(\frac{t}{1/2+\epsilon+\alpha}\right)^{2}}\right)+$
(4.3) $\sum_{1}^{N\rightarrow\infty}\left\{\frac{1}{2n}-\left(\frac{1}{2n+1/2+\epsilon+\alpha}\right)\left(\frac{1}{1+\left(\frac{t}{2n+1/2+\epsilon+\alpha}\right)^{2}}\right)\right\}$
Without Stirling formula as in (A.14 ) , the equivalent relation of $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]$ in (3.22 ) or (3.21), is (4.4). See fig. (2), and fig.(14) for respectively odd and even characters and fig 15 for $\zeta(s)(s-1)$ .
(4.4) $\frac{\partial\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}=\frac{\partial\Im\left\{\ln\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]\right\}}{\partial t}=\frac{1}{2}\ln\left(\frac{q}{\pi}\right)+\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}$
Figure 1. Derivative $\frac{\partial}{\partial\epsilon}$ of (4.4). With $\alpha=0$ even, $\alpha=1$ odd characters. The point of crossing with horizontal axis in case $\alpha=0$ is: $0.585<t_{cross}<0.588$. The curves are independent from congruence modulus $q$.
4.1. Odd characters ($\alpha=1$), $\forall t$, and $\forall q$
In fig 2 is plotted (4.4) For $q=3;\alpha=1$. For $q^{\prime}>q$ we must shift the curve upward. For example, with $q^{\prime}=11$, adding $\frac{\ln(q^{\prime}/\pi)-\ln(3/\pi)}{2}>0.64$. So for $q\geq 11$, (4.4) is positive $\forall t$. Then we are in situation of Theorem 1, because (4.4) is always >0>0, and,
$\frac{\partial}{\partial\epsilon}\frac{\partial\Im\left\{\ln\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]\right\}}{\partial t}=\left\{\frac{\partial}{\partial\epsilon}\frac{\partial}{\partial t}\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]\right\}_{\alpha=1,}>0$ too. See fig. 1 ). So, as in previous section we conclude that for $q\geq 11;\alpha=1$: RH is true $\forall t$. The sufficient condition (4.5)
(4.5) $t:\left[\frac{1}{2}\ln\left(\frac{q}{\pi}\right)+\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}\right]_{\alpha=1}>0\ \ ;\ \ \left\{\frac{\partial}{\partial\epsilon}\frac{\partial}{\partial t}\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]\right\}_{\alpha=1,}>0\$
can be used also locally if the zeros happen at $t$ values in which same situation (4.5) is locally met. The zeros of $L(s,\chi_{primitive})$ distribute like the peaks of the expression (3.13 ) that we know occurs exactly at the zero of $L(s,\chi_{primitive})$. Irrespective if RH is verified or not.
Characters phase $n=0$ $n=1$ $n=2$ $n=3$ $n=4$
$\angle[\chi_{0}(n)]$ undefined 0 0 0 0
$\angle[\chi_{1}(n)]$ undefined 0 $\pi/2$ $(3/2)\pi$ $\pi$
$\angle[\chi_{2}(n)]$ undefined 0 $\pi$ $\pi$ 0
$\angle[\chi_{3}(n)]$ undefined 0 $(3/2)\pi$ $\pi/2$ $\pi$
Table 1. Phase of characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=5$. In column $n=0$ the modulus ( we mean : $\sqrt{\Re^{2}[\chi]+\Im^{2}[\chi]}$ ) is zero. Elsewhere it is 1. When $q$ is prime only principal and primitive characters are present. See [3, p. 168].Notice in each, but the first, row $\sum_{h<q:gcd(h,q)=1}e^{i\ \angle[\chi_{h}]}=0$ for $\chi$ not principal see [3, p. 256]. Table loaded from [11] and adapted for (3.13 ) .
q Odd chrs Suff. Cond. (4.5) closest odd chrs correlation peak distance from $t=0$
33 $|t|>2$ $|t|>8$
44 $|t|>1.5$ $|t|>6$
55 $|t|>1.25$ $|t|>4$
77 $|t|>0.75$ $|t|>2$
88 $|t|>0.5$ $|t|>2.5$
99 $|t|>0.25$ $|t|>2.4$
Table 2. For odd characters, see fig 2, sufficient condition to apply the argument of Theorem 1 pag. 15, is $\left[\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}\right]_{\epsilon=0}+\frac{\ln(q/\pi)}{2}>0$ with $\alpha=1$. Here are reported (second column) his crossing with horizontal axis in fig 2 at various $q^{\prime}$, while in third column the $t$ interval without peaks (zeros). For $q=6$ and $q=10$ no primitive characters are present [15].
Figure 2. Plot of (4.4). For $q^{\prime}>3$ we must shift the curve by the quantity $\frac{\ln(q^{\prime}/\pi)-\ln(3/\pi)}{2}$. For $q\geq 11$ the curve is positive $\forall t$.Notice that the asymptotic result (A.14 ) is well verified at least from $|t|>3$. Here $T_{Asymp}(\alpha)\approx 3$.
For $q<11$ see for example fig. 3 .Here by direct verification, we see that for $q=5$ all zeros are far from negative values of (4.4). We see that the minimum distance of the odd characters ($\chi_{1}$ and $\chi_{3}$) from $t=0$ happen to be $\approx 4>1.25$ where (4.4) with $q=5$ crosses horizontal axis. See also tab. 2. So RH is verified for $q=5$, $\chi_{1}$ and $\chi_{3}$, because at each odd character zero we can locally apply argument of Theorem 1 as sufficient condition (4.5) is locally met. For even primitive character $\chi_{2}$ this is not applicable.
Considering data of tab. 2, built as for case $q=5$, RH is verified for all primitive characters with $3\leq q<10$.
Figure 3. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in not principal character case $q=5$ ( see tab. 1) , with $p_{max}=p^{*}=158\times 10^{6}$ ( here are used the first 8868881 primes in (3.13 ) ).In wolfram demostration project - Dirichlet L-Functions and their zeros we can compute, by zeroing real and imaginary parts, the zeros closer to $t=0$ of what here are called X1 and X3, the result is: $\pm 9.443$, $\pm 8.457$, $\pm 6.184,\ \ \pm 4.133\$. Very comparable with results in figure.
Figure 4. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=5$, with $p_{max}=(p^{*})^{3}=45.8\times 10^{6}$ ( here are used the first 2763823 primes in (3.13 ) ).The peaks are lower and fatter with respect to fig. 3, but, the inter-peaks behavior is much more close to $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$, or, better, to (4.4), that follows different behavior close to $t=0$ if $\alpha=0$ (even chrs) or $\alpha=1$ (odd chrs).
Figure 5. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=3$ and $q=4$, with $p_{max}=(p^{*})^{3}=158\times 10^{6}$ ( here are used the first 8868881 primes).For $q=3$, and, $\chi_{1}(n)=(0,1,-1)$ or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab. 1. For $q=4$, and $\chi_{1}(n)=(0,1,0,-1)$, or, $\angle[\chi_{1}(n)]=(undefined,0,undefined,\pi)$ to comply with format of tab. 1. Notice close to $t=0$ the characteristic behavior of odd primitive L-functions. Compare with fig. 2. Until now the only example of even primitive L-function is phase variation of X2Q5 in fig. 3, and fig. 4. In this case compare with fig. 14.
Putting all pieces together we can state that $\xi(s,\chi_{odd\ primitive})$, is RH compliant $\forall t$ and $\forall q$ .
Characters / n= 0 11 22 33 44 55 66 77 88 99 $10$ $11$ $12$
$\chi_{0}(n)$ 0 1 1 1 1 1 1 1 1 1 1 1 1
$\chi_{1}(n)$ 0 11 $w$ $w^{4}$ $w^{2}$ i-i $w^{5}$ $w^{5}$ $i$ $w^{2}$ $^{4} w-w$ 1-1
$\chi_{2}(n)$ 0 11 $w^{2}$ $w^{2}$ $w^{4}$ 1-1 $w^{4}$ $w^{2}$ 1-1 $w^{4}$ $w^{2}$ $w^{2}$ 11
$\chi_{3}(n)$ 0 11 $i$ 11 1-1 $i$ $i$ i-i i-i 11 1-1 i-i 1-1
$\chi_{4}(n)$ 0 11 $w^{4}$ $w^{4}$ $w^{2}$ 11 $w^{2}$ $w^{2}$ 11 $w^{2}$ $w^{4}$ $w^{4}$ 11
$\chi_{5}(n)$ 0 11 $w^{5}$ $w^{2}$ $w^{4}$ i-i $w$ w-w $i$ $w^{4}$ $w2$ $w^{5}$ 1-1
$\chi_{6}(n)$ 0 11 1-1 11 11 1-1 1-1 1-1 1-1 11 11 1-1 11
$\chi_{7}(n)$ 0 11 w-w $w^{4}$ $w^{2}$ $i$ $w^{5}$ $w^{5}$ i-i $w^{2}$ $^{4} ww$ 1-1
$\chi_{8}(n)$ 0 11 $w^{2}$ $w^{2}$ $w^{4}$ 11 $w^{4}$ $w^{4}$ 11 $w^{4}$ $w^{2}$ $w^{2}$ 11
$\chi_{9}(n)$ 0 11 i-i 11 1-1 i-i i-i $i$ $i$ 11 1-1 $i$ 1-1
$\chi_{10}(n)$ 0 11 $w^{4}$ $w^{4}$ $w^{2}$ 1-1 $w^{2}$ $w^{2}$ 1-1 $w^{2}$ $w^{4}$ $w^{4}$ 11
$\chi_{11}(n)$ 0 11 $w^{5}$ $w^{2}$ $w^{4}$ $i$ w-w $w$ i-i $w^{4}$ $w^{2}$ $w^{5}$ 1-1
Table 3. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=13$, $w=e^{i\pi/6}$. I Table loaded from [11] .
Characters / n= 0 11 22 33 44 55 66 77 88
$\chi_{0}(n)$ 0 1 1 0 1 1 0 1 1
$\chi_{1}(n)$ 0 11 $w$ 0 $w^{2}$ $w^{2}$ 0 w-w 1-1
$\chi_{2}(n)$ 0 11 $w^{2}$ 0 w-w w-w 0 $w^{2}$ 11
$\chi_{3}(n)$ 0 11 1-1 0 11 1-1 0 11 1-1 not primitive
$\chi_{4}(n)$ 0 11 w-w 0 $w^{2}$ $w^{2}$ 0 w-w 11
$\chi_{5}(n)$ 0 11 $w^{2}$ 0 w-w w-w 0 $w^{2}$ 1-1
Table 4. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=9$, $w=e^{i\pi/3}$. I Table loaded from [11] .
Figure 6. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes ).
Figure 7. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab. 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes).
Figure 8. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) flipped vertically. So Lemma 2 is verified.
Figure 9. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes). Notice that inter-peaks level is the same of fig, (14) flipped vertically. So Lemma 2 is verified.
Figure 10. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 11. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 12. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
Figure 13. The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically. So Lemma 2 is verified.
4.2. Even characters ($\alpha=0$)
[…]
In other words, considering also tab. 5, if $L(s,\chi_{primitive})$ does not comply with RH, in $|t|<t_{cross}\approx 0.585$, then, this must occurs necessarily within following constraints.
(4.6) $|t|\leq 0.58\ \ ,\ \ and\ \ ,\ \ 21\leq q\leq 220\ \ ;\ \ \chi_{even\ primitive}$
For $q<21$ no zeros are present in $|t|\leq 0.58$. For $q>220$ Theorm 1 applies in $|t|\leq 0.58$.
q closest even chrs correlation peak distance from $t=0$
55 $|t|>6$
77 $|t|>4$
88 $|t|>4$
99 $|t|>3$
$11$ $|t|>2$
$12$ $|t|>3$
$13$ $|t|>3$
$15$ $|t|>2$
$16$ $|t|>2$
$17$ $|t|>1.5$
1919 $|t|>1.4$
2020 $|t|>2$
Table 5. For even characters, see fig 14, no correlation peaks, i.e, zeros are present in $|t|<t_{cross}\approx 0.58$ at least till $q=20$. For $q=14$ and $q=18$ no primitive characters are present. See also [15].
Figure 14. Plot of (4.4). The conditions of Theorem 1: the curve (4.4) positive, like his $\frac{\partial}{\partial\epsilon}$ derivative, are verified only for $q\geq 220$ and $|t|<t_{cross}$, see fig. 1 . Here $T_{Asymp}(\alpha)\approx 3$.
5. Conclusions
For Theorem 1 in section 3.7, and with extentions in section 4 we can affirm that RH for $\xi((t,\epsilon,\chi_{primitive})$ is true at least for all $L(s,\chi_{odd\ primitive\ })$ q,|t|\ \ $\forall$ q\ ,\ $\forall$|t|. We can say the same of $L(s,\chi_{even\ primitive\ })$ but only within the constraints
$|t|\leq 0.58\ \ ,\ \ and\ \ ,\ \ q>220\ \ ;\ \ \chi_{even\ primitive}$
Application of main idea to $\zeta(s)$ is proposed in a pre-print paper [14].
AcknowledgmentsAcknowledgments
I thank Paolo Lodone for useful discussions, and, contributions.
Appendix A Phase variations of $\xi(s)$, $\xi(s,\chi)$, $\zeta(s)$ and $L(s,\chi)$
Phase variations along t for $\xi(s)$ and $\zeta(s)$ (see (1.2 ) )are connected by:
(A.1) $\left[\frac{\partial\angle[\xi(1/2+\epsilon+it)]}{\partial t}\right]=\frac{\partial\angle[\zeta(s)(s-1)]}{\partial t}+\ln\left[\sqrt{\frac{t}{2\pi}}\right]\quad+O(t^{-2})$
Phase variations along t for $\xi(s,\chi_{primitive})$ and $L(s,\chi_{primitive})$ (see (1.2 )) are connected by:
(A.2) $\left[\frac{\partial\angle[\xi(s,\chi_{primitive}))]}{\partial t}\right]=\frac{\partial\angle[L(s,\chi_{primitive})]}{\partial t}+\ln\left[\sqrt{\frac{tq}{2\pi}}\right]\quad+O(t^{-2})$
( $q$ is the congruence modulus). To prove eq. (A.1), and (A.2) notice that:
(A.3) $\angle\left[\Gamma\left(\frac{s}{2}\right)\frac{s(s-1)}{2}\pi^{-\frac{s}{2}}\zeta(s)\right]=\angle\left[\zeta(s)(s-1)\ \Gamma\left(\frac{s}{2}+1\right)\pi^{-\frac{s}{2}}\right]=\angle\left[\zeta(s)(s-1)\right]+\angle\left[\Gamma\left(\frac{s}{2}+1\right)\pi^{-\frac{s}{2}}\right]$
While for eq. (A.2), and, (1.2 ) we have:
(A.4) $\theta_{primitive}(t)=\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\ \ L(s,\chi)\right]=\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]+\angle\left[L(s,\chi)\right]$
With $\alpha_{1}=\alpha=0$, even $\chi_{primitive}$, or $\alpha_{1}=\alpha=1$, odd $\chi_{primitive}$. See (1.2 ). If we consider the factor (s1)(s-1) attached to $\zeta(s)$ like in (A.1), we can use (A.4) again putting $\alpha_{1}=0\ ,\ \alpha=2\ ,\ q=1$. As $z+1=\frac{s+\alpha}{2}\rightarrow z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$, Stirling formula ( [7, p. 109,112]) allows us to write (A.4) for each of the 3 choices of $(\alpha,\alpha_{1})$:
(A.5) $\ln(\Gamma(z+1))+\ln\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\right]=\ln(\Gamma(z+1))-\frac{s+\alpha_{1}}{2}\ \ln\left(\frac{\pi}{q}\right)=$
$\ln\left(e^{-z}z^{z+\frac{1}{2}}(2\pi)^{\frac{1}{2}}\right)+\left(\sum_{k=1}^{K-1}\frac{B_{2k}}{2k(2k-1)z^{2k-1}}\right)+R_{2K}(z)-\ln\left(\frac{\pi}{q}\right)\frac{s+\alpha_{1}}{2}=\Re_{1}+\Re_{2}+\Re_{3}+i(\Im_{1}+\Im_{2}+\Im_{3})$
Although the expression (A.5) is a not-convergent asymptotic expansion, it can be used with a finite $K$ keeping the modulus of error term $|R_{2K(z)}|$ to a suitable level. Let us call $\Im_{1}$ the imaginary of asymptotic (for $t$ big) part of $\ln\left(e^{-z}z^{z+\frac{1}{2}}(2\pi)^{\frac{1}{2}}\right)$. And $\Re_{1}$ is the real part. Let us call $\Im_{2}$ the left imaginary part of $\ln\left(e^{-z}z^{z+\frac{1}{2}}(2\pi)^{\frac{1}{2}}\right)$ wich goes to zero for $t\rightarrow\infty$. And $\Re_{2}$ is the real part.And finally let us call $\Re_{3}+i\Im_{3}=\left(\sum_{k=1}^{K-1}\frac{B_{2k}}{2k(2k-1)z^{2k-1}}\right)+R_{2K}(z)$.
The $B_{j}$ are the Bernoulli numbers ( [7, p. 11] ), that vanish for odd $j$ while:
$B_{2}=\frac{1}{6}\quad;\quad B_{4}=-\frac{1}{30}\quad;\quad B_{6}=\frac{1}{42}\quad;\quad B_{8}=-\frac{1}{30}\quad;\quad B_{10}=\frac{5}{66}\quad;\quad B_{12}=-\frac{691}{2730}\quad$...
The modulus of the error term $|R_{2K}(z)|$ is bounded by:
(A.6) $|R_{2K}(z)|<\left(\frac{B_{2K}}{2K(2k-1)z^{2K-1}}\right)\frac{1}{\left[\cos\left(\frac{\arg(z)}{2}\right)\right]^{2K}}$
where $\arg(z)$ is taken in the interval: $\pi<\arg(z)<\pi$ (see [7, p. 112].
In (A.5) we may put : $z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$ , so, leaving apart only $\Im_{3}$ for now, we have:
$\ln\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]\approx\Re_{1}+\Re_{2}+i(\Im_{1}+\Im_{2})=$
$\left(\frac{s+\alpha}{2}-1\right)+\left(\frac{s+\alpha}{2}-\frac{1}{2}\right)\ln\left(\frac{s+\alpha}{2}-1\right)+\ln(\sqrt{2\pi})-\ln\left(\frac{\pi}{q}\right)\frac{s+\alpha_{1}}{2}=$
$\frac{3}{4}-\frac{\epsilon+\alpha}{2}-\frac{it}{2}+\left(\frac{\epsilon+\alpha}{2}-\frac{1}{4}+\frac{it}{2}\right)\left\{\ln\sqrt{\left(\frac{\epsilon+\alpha}{2}-\frac{3}{4}\right)^{2}+\frac{t^{2}}{4}}+i\arctan\left[\Im=\frac{t}{2};\Re=\frac{\epsilon+\alpha}{2}-\frac{3}{4}\right]\right\}+$
(A.7) $\frac{\ln(2\pi)}{2}-\ln\left(\frac{\pi}{q}\right)\left(\frac{\epsilon+\alpha_{1}}{2}+\frac{1}{4}\right)+i\frac{t}{2}\ln\left(\frac{q}{\pi}\right)$
Taking into account that: $\arctan\left[\Im=\frac{t}{2};\Re=\frac{\epsilon+\alpha}{2}-\frac{3}{4}\right]=\frac{\pi}{2}+\arctan\left[\frac{3-2(\epsilon+\alpha)}{2t}\right]$, we get:
(A.8) $\frac{3}{4}-\frac{\epsilon+\alpha}{2}-\frac{it}{2}+\left(\frac{\epsilon+\alpha}{2}-\frac{1}{4}+\frac{it}{2}\right)\left\{\ln\sqrt{\left(\frac{\epsilon+\alpha}{2}-\frac{3}{4}\right)^{2}+\frac{t^{2}}{4}}+i\left\{\frac{\pi}{2}+\arctan\left[\frac{3-2(\epsilon+\alpha)}{2t}\right]\right\}\right\}+$
$\frac{\ln(2\pi)}{2}-\ln\left(\frac{\pi}{q}\right)\left(\frac{\epsilon+\alpha_{1}}{2}+\frac{1}{4}\right)+i\frac{t}{2}\ln\left(\frac{q}{\pi}\right)$
Let us choose $T_{Asymp}(\alpha)$, in order for $t>T_{Asymp}(\alpha)$ the braces in (A.8) becomes $\left\{\ln\left(\frac{t}{2}\right)+i\frac{\pi}{2}\right\}$,so we have:
$\ln\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]\approx\Re_{1}+i\Im_{1}=$
$\frac{3}{4}-\frac{\epsilon+\alpha}{2}-\frac{it}{2}+\left(\frac{\epsilon+\alpha}{2}-\frac{1}{4}+\frac{it}{2}\right)\left\{\ln\left(\frac{t}{2}\right)+i\frac{\pi}{2}\right\}+\frac{\ln(2\pi)}{2}-\ln\left(\frac{\pi}{q}\right)\left(\frac{\epsilon+\alpha_{1}}{2}+\frac{1}{4}\right)+i\frac{t}{2}\ln\left(\frac{q}{\pi}\right)$
From which we find the the asymptotic imaginary parts $\Im_{1}$ :
(A.9) $\Im_{1}=\theta_{primitive}(t)=-\frac{t}{2}+\frac{t}{2}\ln\left(\frac{tq}{2\pi}\right)-\frac{\pi}{8}+\frac{\pi}{4}(\epsilon+\alpha)$
From (A.8) the imaginary part $\Im_{2}$ that goes to zero as $t\rightarrow\infty$:
(A.10) $\Im_{2}:=\frac{t}{2}\ln\sqrt{1+\left(\frac{2(\epsilon+\alpha)-3}{2t}\right)^{2}}+\left(\frac{\epsilon+\alpha}{2}-\frac{1}{4}\right)\arctan\left(\frac{3-2(\epsilon+\alpha)}{2t}\right)$
for $t$ big, posing $y=\epsilon+\alpha$:
(A.11) $\Im_{2}\approx\frac{t}{4}\left(\frac{2(\epsilon+\alpha)-3}{2t}\right)^{2}+\left(\frac{\epsilon+\alpha}{2}-\frac{1}{4}\right)\left(\frac{3-2(\epsilon+\alpha)}{2t}\right)=\frac{t}{4}\left(\frac{2(y-3)}{2t}\right)^{2}+\left(\frac{y}{2}-\frac{1}{4}\right)\left(\frac{3-2y}{2t}\right)$
$\frac{\partial^{2}\Im_{2}}{\partial y\partial t}=-\frac{4(2y-3)}{16t^{2}}-\frac{1}{2}\frac{3-2y}{2t^{2}}+\left(\frac{y}{2}-\frac{1}{4}\right)\frac{2}{2t^{2}}$
For $\epsilon=0$ and $\alpha=2$ we have:
$\left[\frac{\partial^{2}\Im_{2}}{\partial\epsilon\partial t}\right]_{\epsilon=0\ ,\ \alpha=2,\ \alpha_{1}=0}=-\frac{4}{16t^{2}}+\frac{1}{4t^{2}}+\frac{3}{4}\frac{2}{2t^{2}}=\frac{3}{4t^{2}}$
For $\epsilon=0$ and $\alpha=1$ we have:
$\left[\frac{\partial^{2}\Im_{2}}{\partial\epsilon\partial t}\right]_{\epsilon=0\ ,\ \alpha_{1}=\alpha=1}=+\frac{4}{16t^{2}}-\frac{1}{4t^{2}}+\frac{1}{4}\frac{2}{2t^{2}}=\frac{1}{4t^{2}}$
For $\epsilon=0$ and $\alpha=0$ we have:
$\left[\frac{\partial^{2}\Im_{2}}{\partial\epsilon\partial t}\right]_{\epsilon=0\ ,\ \alpha_{1}=\alpha=0}=+\frac{12}{16t^{2}}-\frac{3}{4t^{2}}-\frac{1}{4}\frac{2}{2t^{2}}=-\frac{1}{4t^{2}}$
While $\Im_{3}=\Im\left[\left(\sum_{k=1}^{K-1}\frac{B_{2k}}{2k(2k-1)z^{2k-1}}\right)+R_{2K}(z)\right]$ for $K=3$.So k=1,and 2k=1,and\ 2
(A.12) $\Im_{3}-R_{2K}(z)=\frac{-1}{6t\left[1+\left(\frac{2\epsilon+2\alpha-3}{2t}\right)^{2}\right]}-\frac{1}{45t^{3}\left[1+\left(\frac{2\epsilon+2\alpha-3}{2t}\right)^{2}\right]^{3}}+\frac{(2\epsilon+2\alpha-3)^{2}}{60t^{5}\left[1+\left(\frac{2\epsilon+2\alpha-3}{2t}\right)^{2}\right]^{3}}$
[…]
Besides : $\frac{\partial\Im_{3}}{\partial t}\approx\frac{1}{6t^{2}}+3\frac{(2\epsilon+2\alpha-3)^{2}}{24t^{4}}$. While $\frac{\partial^{2}\Im_{3}}{\partial\epsilon\partial t}\approx\frac{(2\epsilon+2\alpha-3)}{4t^{4}}$, so, the change with $\epsilon$ is negligible. The derivative of (A.9) with respect to $t$ is:
(A.13) $\frac{\partial\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha_{1}}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}=\frac{\partial\Im_{1}}{\partial t}+\frac{\partial\Im_{2}}{\partial t}+\frac{\partial\Im_{3}}{\partial t}$
where:
(A.14) $\frac{\partial\Im_{1}}{\partial t}=\ln\left(\sqrt{\frac{tq}{2\pi}}\right)\quad;\quad q=1\ for\ \zeta(s)\ ;\ q>1\ for\ L(s,\chi_{primitive})\ ;\ t>T_{Asymp}(\alpha)$
Notice that $T_{Asymp}(\alpha)$, allows us to use the asymptotic relation (A.14 ). Following (A.4) it cannot depend on $q$ or on $\alpha_{1}$, but only on $\Gamma\left(\frac{s+\alpha}{2}\right)$.
$\frac{\partial^{2}\Im_{3}}{\partial\epsilon\partial t}$ is negligible, and $\frac{\partial}{\partial\epsilon}\frac{\partial\Im_{1}}{\partial t}=0$. So doing $\frac{\partial}{\partial\epsilon}$ of (A.13) we have the only three cases of $\frac{\partial^{2}\Im_{2}}{\partial\epsilon\partial t}$ ( from (A.13 ) :
(A.15) $\overbrace{\left[\frac{\partial}{\partial\epsilon}\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}\right]_{\epsilon=0}=\frac{3}{4t^{2}}}^{\zeta(s)\ :\alpha=2}\ \ ;\ \overbrace{\left[\frac{\partial}{\partial\epsilon}\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}\right]_{\epsilon=0}=\frac{1}{4t^{2}}}^{\chi_{primitive}(-1)=-1\ :\alpha=1}\ \ ;\ \ \overbrace{\left[\frac{\partial}{\partial\epsilon}\frac{\partial\angle\left[\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}\right]_{\epsilon=0}=-\frac{1}{4t^{2}}}^{\chi_{primitive}(-1)=1\ :\alpha=0}$
Figure 15. Plot of (4.4) for $\alpha=2$ and $q=1$, i.e. for the $\frac{\partial\angle[\zeta(s)(s-1)]}{\partial t}$ function. See (A.1). Notice that the asymptotic phase variation with $t$ (A.14 ) is the same surely for $t>10$, i.e. $T_{Asymp}(\alpha)\approx 10$ here.
Appendix B differences between (3.14) and (3.15)
The difference between the infinite sum (3.13) and the same infinite sum with substitution of (3.14) with (3.15) is always a finite number at least for $\epsilon\geq 0$. So the correlation peaks of (3.13) stemming from the zeros of $L(s,\chi)$, see fig. 3, are present also in (3.13) after subsitution.
PROOF. Using Taylor formula for arctangent the difference between (3.13) using (3.14) or (3.15) can be written as:
$\left.\frac{\ln(p^{*})}{2\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right|_{t_{1}}^{t_{2}}+$
$\frac{\ln(p^{*})}{2\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\frac{2\cos[\ln(p)t-\angle[\chi(p)]]\sin\left(\frac{\ln(p)\pi}{\ln(p^{*})}\right)}{p^{1/2+\epsilon}}=$
(B.1) $\left.\frac{\ln(p^{*})}{2\pi}\ \ \ \sum_{p<p_{max}\ :\ gcd(p,q)=1}\ \ \ \ \ \sum_{n\ odd>1}\frac{(-1)^{(n-1)/2}}{n}\left[\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)^{n}\right]\right|_{t_{1}}^{t_{2}}-$
(B.2) $\left.\frac{\ln(p^{*})}{2\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])\cos(\ln(p)t-\angle[\chi(p)])}{[p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])]p^{1/2+\epsilon}}\right)\right|_{t_{1}}^{t_{2}}$
Developing last sum we have, focusing on main term:
(B.3) $\left.\frac{\ln(p^{*})}{4\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\left(\frac{\sin(2(\ln(p)t-\angle[\chi(p)]))}{[p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])]p^{1/2+\epsilon}}\right)\right|_{t_{1}}^{t_{2}}\approx$
(B.4) $\frac{\ln(p^{*})}{4\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\left(\frac{\sin(2(\ln(p)t_{2}-\angle[\chi(p)]))-\sin(2(\ln(p)t_{1}-\angle[\chi(p)]))}{[p^{1/2+\epsilon}-\cos(\ln(p)t_{2}-\angle[\chi(p)])][p^{1/2+\epsilon}-\cos(\ln(p)t_{1}-\angle[\chi(p)])]}\right)$
It is apparent that the difference between computation (3.13 ), upon substitution of (3.14) with (3.15) is absolutely convergent for sure from $\epsilon>0$ and conditionally convergent for $\epsilon=0$.
Proof. Let us look at (3.27), and, consider only one $h$-class primes till $p_{max}$ ( as in (3.18 ).
Each equal sign interval (see (3.27)) now $\Delta p=\frac{\pi}{2t}p$ alternating sign, in (B.4) is divided by $\approx p$. But the prime density is $\frac{1}{\ln(p)}$. So we have an alternating sign decreasing module series. The limit must be finite by Leibnitz rule[1, p. 404]. So the divergences at $L(s,\chi)$ zeros, i.e. correlation peaks, are the same.
The double sum, in $p$ and in nodd>1n\ odd>1, as the neglected term between (B.3) and (B.4) cannot influence divergences all the same, because of the common factor $\sim p^{-3(1/2+\epsilon)}$.
If we consider all the $\phi(q)$ terms ( possible if $p_{max}<\infty$), even if we cannot conclude that the sum of all $\phi(q)$ contributions tend to zero ( as in Lemma 4), we can presume that the poor convergence of a single $h$-class primes is enanced when all $h$-classes are combined together. This seems confirmed by figg. 17, 18, and, 19.
Figure 16. Comparative plot of (3.13) using (3.14) or (3.15), with $t$ fixed, $\angle[\chi]=0$, and continuously varying pp. It is reported also (3.13) with substitution of intermediate expression $\left\{\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}$
Figure 17. Comparative plot of (3.13) computed with approximation (3.15), with $\epsilon=0$, and $390<t<395$, for $\chi_{1}$ and $q=5$. See tab. 1. We have used the first 2.763.823 primes like in following figures.
Figure 18. Comparative plot of (3.13) computed with approximation (3.15), with $\epsilon=0$, and $900<t<905$, for $q=3$, and, $\chi_{1}(n)=(0,1,-1)$, to comply with format of tab. 3 . Or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab. 1.
Figure 19. Comparative plot of (3.13) computed with approximation (3.15), with $\epsilon=0$, and $900<t<905$, for $q=4$, and $\chi_{1}(n)=(0,1,0,-1)$, to comply with format of tab. 3. Or, $\angle[\chi_{1}(n)]=(undefined,0,undefined,\pi)$ to comply with format of tab. 1. Notice in interval $904<t<904.5$ there are two zeros not resolved by the implicit resolution: $\frac{2\pi}{\ln(45.8\times 10^{6})}\approx 0.356$
Appendix C Convergence for primitive characters in critical strip
Let us see a way to show that
(C.1) $\L (s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}$
converge in any compact on critical strip.
Abel summation formula , in the version of [19], is reported for easy reading. Let $\{a_{n}\}_{n=0}^{\infty}$ be a sequence of real or complex numbers. Define partial sum function A by $A(t)=\sum_{0\leq n\leq t}a_{n}$ for any real number $t$. Fix real number $x<y$ and let $\phi$ a continuously differentiable function on $[x,y]$. Then $\sum_{x<n\leq y}a_{n}\phi(n)=A(y)\phi(y)-A(x)\phi(x)-\int_{x}^{y}A(u)\phi^{\prime}(u)du$
Posing
(C.2) $K(X)=\sum_{0<x\leq X}\chi(x)$
we have:
(C.3) $\sum_{n=1}^{n=N=\lfloor X\rfloor}\chi(n)\times\left(\frac{1}{n^{s}}\right)=K(X)\times\left(\frac{1}{X^{s}}\right)-K(1)\times\left(\frac{1}{(1)^{s}}\right)-(-s)\int_{1}^{X}\frac{K(u)du}{u^{s+1}}$
as for $X\rightarrow\infty$:
  • $K(X)$ is bounded. For sure it is $\phi(q)$.
  • $K(X)\times\left(\frac{1}{X^{s}}\right)\rightarrow 0$, for $\Re(s)>0$
  • $K(1)\times\left(\frac{1}{1^{s}}\right)=1$
  • $\int_{1}^{X}\frac{K(u)du}{u^{s+1}}$ converges for $\Re(s)>0$ because of first bullet.
So $\L (s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}$ converges in $\Re(s)>0$
The crucial point for series (C.1) is the bounded behavior of (C.2).
Appendix D Some feedback from readers
I thank all the readers who have send to me comments on previous versions of the paper. Below is given a list of them (unified following subject) with my answers. They are useful especially for me to delve deeper in this hard subject.
  • Invalid Use of the Euler Product in the Critical Strip. If it is correct that the Dirichlet series $\L (s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}$ converges for $\Re(s)>0$ (for non-principal character) and that the L-function has an Euler product. However, that these two facts imply the equality $\sum_{n}=\prod_{p}$is incorrect for $\Re(s)>0$. The proof of the identity, first discovered by Euler, relies on multiplying out the infinite geometric series for each prime factor and rearranging the order of an infinite number of terms. This rearrangement is only guaranteed to preserve the value of the sum if the series is absolutely convergent. So section 3 is baseless. .
    ANSW.I agree on the necessity to fill this gap (not a mistake I think) in the logical path of the paper. In new version of section 3 I have shown that Euler product makes sense also in critical strip for primitive $L$-Functions. Equation (1.1) does not hold for $0<\Re(s)<1$, but, subsection 3.1 shows a way in order to exploit Euler product for variations, in particular phase variations, on small $t$-intervals. On the other hand ( even if it is not an argument) it seems to me worth to draw the attention on figg, 3 through 13 where it is apparent that (3.9), at least numerically, works rather well even with only $m=3$. In conclusion it seems to me that subsection 3.1 is a valid base for the developments of section 3. Anyhow I thank the reader for this observation that prompted me to delve deeper on subject.
  • Incorrect Manipulation of Conditionally Convergent Series. The argument in Lemma 6 is invalid. The series for the phase is conditionally convergent, not absolutely convergent. In a conditionally convergent series, the order of summation matters
    ANSW. I agree with the statement, but it is not applicable.The “manipulation” of terms in (3.34 ) (or in others occasions (3.1), (3.18) and (3.20) ) is applied only at addends with index less than the running index. So on a set of finite numbers. Partial sums do not change (as it is stressed thoughout). To explain better. Consider an example $S_{j}=-1+1/2-1/3+1/4-.....=\sum_{j}(-1)^{j}/(j)\rightarrow-\ln(2)\ ,$ as $j\rightarrow\infty$.
    If we take the sum till $J_{max}$ and then we partition the addends with $j\leq J_{max}$ in positive $\sum^{+}_{J_{max}}$ and negatives $\sum^{-}_{J_{max}}$, then we can write $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$, and, we do not alter final result $\ln(2)$ for $J_{max}\rightarrow\infty$ transforming the limit in something like $\infty$-$\infty$. It is exactly the “manipulation” in expression : (3.34 ). While in (3.1), in (3.18), and, in (3.20), we manipulate to highlight some structure of the series, but , with indexes less than the running index . A completely rearranging of terms as in [1, p. 411] without change in final result is possible only where convergence is absolute. But, it is easy to see that there are infinite rearrangement of terms that preserve the sum $\ln(2)$ of example above. For example if we change odd with even terms, i.e. $S^{\prime}_{j}=-1/2+1-1/4+1/3-1/6+1/5-1/8+1/7-.....=-\ln(2)$.
    Partial sums are the same for even index of the new series and different for odd index of new series. So partial sums of $S^{\prime}_{j}$ and $S_{j}$ are different, but they tend always to same limit as $J_{max}\rightarrow\infty$. Notice however we have done a reordering on infinite terms. So we even can say: not all , but, some reordering of infinite terms can be done also on conditional convergent series and the result when index $\rightarrow\infty$ is preserved. But here we do only “manipulations” with indexes less than the running index . So observation does not apply.
  • The argument in equations ( 3.43 ) through (3.50) rests on the idea that because each term in the series for $\epsilon>0$ is smaller in magnitude than the corresponding term for $\epsilon=0$ (due to the factor of $p^{-\epsilon}$), the overall sum must also be smaller in magnitude. This reasoning is incorrect.
    ANSW. ( 3.43 ) is a ratio not of the whole sum, but only of the positive contributions at different $\epsilon$ values. Each positive contribution decreases with increasing $\epsilon$, the same holds for their sum, so this ratio ( 3.43 ) also $\forall p_{max}<\infty$ is less than 1 and also (3.45 ) are completely justyfied. The same holds for negative contributions. Overall sum is not in question at this stage. The factor $p^{-\epsilon}$ is cited to give an idea. The computation takes in consideration correct contributions.
  • Analyzing the sum of positive terms in isolation from the sum of negative (like in (3.45 ) terms is not a valid operation
    ANSW Generally speaking I can agree but (see example two bullets above) $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$ tends exacly to the limit. So I do not understand the sentence: it“is not a valid operation”.
  • algebraic manipulations from ( 3.43 ) through (3.50) effectively treat the series as if it were a sum of positive terms, ignoring the crucial role of cancellation. This is a classic error when dealing with conditional convergence
    ANSW.We have chosen to compute the single pieces of equi-sign contribution in absolute value and to put correct sign given by (3.33) afterward. This is a choice of opportunity because we are interested in a form $\infty$-$\infty$ that, as explained three bullets above, can be unusual, but, it is correct because we consider always indexes less than the running index .
  • It is true that one can make rearrangements when there are finitely many terms, but then one cannot take the infinite limit because rearrangements are not correct for infinite sums. Therefore, the argument in Lemma 6 and the conclusions drawn from it are invalid. ANSW. Perhaps there is a misunderstanding because (see example four bullets above) If we write $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$, we do not alter partial sum and so final result, as we use indexes less than the running index . We transfom the limit in something like $\infty$-$\infty$ without changing result.
  • It seems that in Lemma 4 and the subsequent analysis in Lemma 6,sums over prime numbers are replaced with integrals weighted by the logarithmic integral, citing the Prime Number Theorem in Arithmetic Progressions (PNTAP) in equation (3.7). The PNTAP provides an asymptotic relationship. It states that $\pi(x)_{h,q}\sim Li(x)/\phi(q)$ as $x\rightarrow\infty$. This does not mean the two are equal, and the error term is not negligible. Using this as a direct substitution in an argument that depends on the precise cancellation of oscillating terms is not rigorous. A valid argument would require the use of an explicit formula for the distribution of primes, complete with error terms, and demonstrate that these error terms do not affect the final result.
    ANSW In Lemma 4 it is used the function $Li(x)$, but, with whatsoever well behaved function the result of (3.19 ) is null because it depends from (3.17). So in Lemma 4 there is no reference to (PNTAP). The rationale of (3.20) is that of an algebraic trick to insert, with zero effective effect, the integral of PNT in (3.13). The difference in (3.20) means nothing because we are subtracting zero (i.e.(3.19 )) from (3.13). But it is useful in Lemma 6 to build not differences between primes distribution $\pi(x)$ and $Li(x)$ but ratios between functions involving $\pi(x)$ and $Li(x)$. In ratios asymptotic relations can be used at best. Of course not in differences. So I recommend a more careful reading and I think that the correct expectations of the reader will be perhaps satisfied.
    I thank anyway a lot the reader for this observation that highlight a basic and simple idea. Namely that it is more convenient to look for ratios between functions involving $\pi(x)$ and $Li(x)$ rather than to look for differences.
  • The proposed methods are unconventional ANSW Perhaps you are right, but, my problem is to identify, and, if possible eliminate, eventual flaws.
  • The paper has not be validated by a rigorous peer review ANSW That is true, but before to ask for it ( that by the way it is no simple matter) I hope to uncover possible problems that could invalidate the paper, and at same time I try to stabilize the treatment of the topic. To this aim readers observations are very important for me. After all, as far as I know, these ideas have “circulated” only since one year.
  • Computed results cannot be taken as proofs ANSW Yes I agree. Let us say they are a necessary condition not a sufficient one. However I think that the matching of the $t_{peak}$ of the peaks of (3.13) with the zeros of $L(s,\chi)$, and the matching of the floor with the computed one by (A.14 ) and (4.4) (see on figg, 3 through 13) is an interesting fact.
  • The attempt to link mathematical problems to other fields like physics is not inherently frivolous, but, the lack of peer-reviewed validation makes it a speculative effort rather than a serious approach ANSW The reference to a kinematic quantity as angular momentum is only a short way to describe the method I explored. So I used it in the abstract of [14] only to be synthetic. The importance of this formal link (however mathematically based) with classical mechanics in the logic of the paper is zero. Relating peer-review, I’ll be glad to face it. See note two bullets above.
  • The term “spectrum of primes” is not a rigorously defined concept ANSW Perhaps you are right. I borrowed in ([14]) this term by [16]. But anyhow a formula that gather zeros of L-functions and primes ( as (3.13) ), seems interesting.

### v16

@@
 1. Introduction
-Using Dirichlet characters $\chi(n)$, L functions, [3, p. 249, and 262], and, specially, ([6, p. 3, and 37] (whose conventions, i.e. $\chi_{0}$ as principal, is used throughout) are defined as :
+Using Dirichlet characters $\chi(n)$, L functions, [3, p. 249, and 262], and, specially, ([7, p. 3, and 37] (whose conventions, i.e. $\chi_{0}$ as principal, is used throughout) are defined as :
 (1.1) $\L (s.\chi)=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}=\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad,\quad\quad\quad\Re(s)>1$
@@
 168 -170].
-Companion function like $\xi(s)$, ( [6, p.
-62] ) for $\zeta(s)$, is $\xi(s,\chi_{primitive})$, ( [6, p.
+Companion function like $\xi(s)$, ( [7, p.
+62] ) for $\zeta(s)$, is $\xi(s,\chi_{primitive})$, ( [7, p.
 71] ) for $L(s,\chi_{primitive})$, below their definition :
@@
 While $\alpha=0$ if $\chi_{primitive}(-1)=1$, even character, and, $\alpha=1$ if $\chi_{primitive}(-1)=-1$, odd character.
-Both (1.2 ) have the same zeros of the $\zeta(s)$ or $L(s,\chi)$ functions respectively in the critical strip [6] [7, p.
+Both (1.2 ) have the same zeros of the $\zeta(s)$ or $L(s,\chi)$ functions respectively in the critical strip [7] [8, p.
 16].
-Following [6, p.
+Following [7, p.
 71, and 79] the functional equation for $\xi(s)$ and $\xi(s,\chi)$ are:
 (1.3) $\quad\xi(1-s)=\xi(s)\ \ \ ;\ \ \ \xi(1-s,\bar{\chi})=\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\xi(s,\chi)\quad with\quad\chi_{primitive}\ \ \quad$
-Where $\bar{\chi}$ is the complex conjugate of $\chi$ and $\tau(\chi)$ is the gauss sum $\tau(\chi):=\sum_{m=1}^{q}\chi(m)e^{2\pi im/q}$ see [6, p. 65] , or [3, p. 165].
+Where $\bar{\chi}$ is the complex conjugate of $\chi$ and $\tau(\chi)$ is the gauss sum $\tau(\chi):=\sum_{m=1}^{q}\chi(m)e^{2\pi im/q}$ see [7, p. 65] , or [3, p. 165].
 While zeros of $\zeta(s)$ have also a symmetry with respect to real axis, the $L(s,\chi)$ functions maintain this symmetry only for real characters $\chi$.
@@
 The symbol $\angle[a]$ means phase of complex number $a$.
-  • In appendix B we compare an useful approximation to (3.13 ).
+  • In appendix B we compare an useful approximation to (3.16 ).
   • In appendix C is justified the use of $L(s,\chi_{primitive})$ in critical strip.
@@
 Let us apply Lemma 1 to 1.3 ( remember $\chi_{primitive}$).
-For [6, p.
+For [7, p.
 66] $|\tau(\chi)|^{2}=q$, then $\left|\frac{i^{\alpha}q^{\frac{1}{2}}}{\tau(\chi)}\right|=1$.
 But, at $\epsilon=0$ we have: $1-s=\bar{s}$.
-Besides, as $\Gamma(\bar{s})=\int_{0}^{\infty}x^{\bar{s}-1}e^{-x}dx=\bar{\Gamma}(s)$ ([7, p.
+Besides, as $\Gamma(\bar{s})=\int_{0}^{\infty}x^{\bar{s}-1}e^{-x}dx=\bar{\Gamma}(s)$ ([8, p.
 8]), $\left(\frac{q}{\pi}\right)^{\frac{\bar{s}+\alpha}{2}}=\bar{\left(\frac{q}{\pi}\right)^{\frac{s+\alpha}{2}}}$, and $L(\bar{s},\bar{\chi})=\bar{L}(s,\chi)$ ( (1.1)) , then for 1.2 ), and ( 1.3):
@@
 255] ), and, (1.1) converges absolutely for $\Re(s)>1$, while converges conditionally (see [1, p.
-406] ) for $\Re(s)>0$ ( [10, p.
+406] ) for $\Re(s)>0$ ( [11, p.
 7], and, also unifomily [5, p.
@@
 (3.2) $N^{o}(terms)=3\prod_{j=1}^{j_{max}-1}\left(\left\lceil\frac{\ln(p_{max})}{\ln(p_{j})}\right\rceil+1\right)\ \ ;\ \ last\ factor\left\lceil\frac{\ln(p_{max})}{\ln(p_{max})}\right\rceil+1\ \ ;\ \ is\ put\ to\ 3$
-For example for $p_{max}=31$ we have in (3.2 ) that $N^{o}(terms)=787320$, whoose only 31 belong to $L(s,\chi,p_{max})$. The others belong to $R(s,\chi,p_{max})=R(t,\epsilon,\chi,p_{max})$.
-So we does not need to care, in applying Euler rearrangement for $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, about absolute convergence of the $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime$, because we have an huge, but, finite amount of terms.
+For example for $p_{max}=31$ we have in (3.2 ) that $N^{o}(terms)=787320$, whoose only 31 belong to $L(s,\chi,p_{max})$.
+The others belong to $R(s,\chi,p_{max})=R(t,\epsilon,\chi,p_{max})$.
+While $\rho_{max}(p_{max})$, i.e.
+the greatest $\rho$ in (3.1), is $Primorial(p_{max})$, so an huge number too, though finite.
+Besides if we consider a successions of
+(3.3) $p_{max}^{k}=\rho_{max}^{k-1}(p_{max}^{k-1})$
+it is apparent that all the terms of $R(t,\epsilon,\chi,p_{max}^{k-1})$, are absent in $R(t,\epsilon,\chi,p_{max}^{k})$.
+So it is apparent that a build-up toward a certain value, from whatever $p^{\prime}_{max}$ to $p^{\prime\prime}_{max}\rightarrow\infty$, it is to be excluded for $R(t,\epsilon,\chi,p^{\prime\prime}_{max})$ when $p^{\prime\prime}_{max}\geq\rho_{max}(p^{\prime}_{max})$.
+And we are speaking always of huges but finite set of terms.
+So we does not need to care, in applying Euler rearrangement for $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, and $R(s,\chi,p_{max})$, about absolute convergence of the $\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime$, because we are dealing with huges, but, finite amount of terms.
 We have:
@@
 $N^{o}$ of addends of $R(t,\epsilon,\chi,p_{max})$ is : $3\prod_{j=1}^{j_{max}-1}\left(\left\lceil\frac{\ln(p_{max})}{\ln(p_{j})}\right\rceil+1\right)\ \ -p_{max}$.
+It is convenient to think the finite sum in (3.4) as ordered by growing $\rho$ values. We have $p_{max}<\rho<\rho_{max}(p_{max})$. The latter is bigger than $Primorial(p_{max})$, but finite.
 Notice that variation in $\Delta t$ is due almost completely to $\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]$, because with $p_{j}<<\rho$ we have $\ln(p_{j})<<\ln(\rho)$, and for big $p_{j}\leq p_{max}$, for $\epsilon>0$, we have $|\chi(p_{j})p_{j}^{-s})^{-1}|<<1$ for $p_{j}$ big enough.
 Besides we are interested in $\Delta t$ small.
-Only first term in right part of (3.1), i.e $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, tend to (C.1), so, due to (3.3), we have, considering the whole critical strip:
+Only first term in right part of (3.1), i.e $L(s,\chi,p_{max})=\sum_{n=1}^{p_{max}}\frac{\chi(n)}{n^{s}}$, tend to (C.1). So, considering the whole critical strip, we have:
 (3.5) $\L (s.\chi_{primitive})=\sum_{n=1}^{\infty}\ \chi(n)/n^{s}\neq EP(t,\epsilon,p_{max}\rightarrow\infty)=\prod_{\forall p}\frac{1}{1-\frac{\chi(p)}{p^{s}}}\quad p\quad prime\quad,\quad\quad\quad 0<\Re(s)<1\ \ \$
@@
 Notice however that our aim is to use $\angle[L(s,\chi)]$, and, also the phase is undefined when $L(s,\chi)=0$.
-But if we consider the difference (or better the phase difference) in a finite interval, suitably far from $\L (s.\chi)=0$, the conclusion is different. Let us take:
+If we consider the difference (or better the phase difference) in a finite interval, suitably far from $\L (s.\chi)=0$, the conclusion is different. Let us take:
 […]
@@
 (3.7) $\Delta\left[R(t,\epsilon,\chi,p_{max})\right]_{\Delta t}=\int_{t_{1}}^{t_{2}}\frac{\partial R(t,\epsilon,\chi,p_{max})}{\partial t}dt$
-where in (3.3) we have an huge but finite number of terms with zero mean value.
-If $p_{max}>>p^{*}$ we have a lot of cancellation in $\Delta t$ interval for single (3.6) addend.
+where:
+(3.8) $\frac{\partial R(t,\epsilon,\chi,p_{max})}{\partial t}=\sum_{\rho>p_{max}}-\ln(\rho)[\sin(\ln(\rho)t)+i\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}\prod_{\alpha_{j}>0}(1-\chi(p_{j})p_{j}^{-s})^{-1}+$
+$\sum_{\rho>p_{max}}[\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}\times$
+$\left\{\sum_{j\ ,\ \alpha(j)\neq 0}-\ln(p_{j})\frac{[\sin(\ln(p_{j})t)+i\cos(\ln(p_{j})t)]}{1-\chi(p_{j})p_{j}^{-s})^{2}}\frac{\chi(p_{j})}{p_{j}^{\Re[s]}}\prod_{\alpha_{k}>0\ k\neq j}(1-\chi(p_{k})p_{k}^{-s})^{-1}\right\}$
+In (3.8) we have an huge but finite number of terms with zero mean value ad with phase variation given subtantially by $\ln(\rho)t$.
+If $p_{max}>>p^{*}$ we have a lot of cancellation in $\Delta t$ integral (3.7) for single $\rho$ (3.8)addend.
 The cancellation will be complete, for a particular $\rho$, only if
-(3.7) $\Delta t\times\ln(\rho)=\frac{2\pi}{\ln(p^{*})}\ln(\rho)=m\times 2\pi\ ;\ m\in N\ ,\ i.e.\ \rho=(p^{*})^{m}$
+(3.9) $\Delta t\times\ln(\rho)=\frac{2\pi}{\ln(p^{*})}\ln(\rho)=\mu\times 2\pi\ ;\ with\ \mu=m\in N\ ,\ i.e.\ \rho=(p^{*})^{m}$
 When condition (3.9) is not verified the derivative of a term of (3.1, in (3.7) integral, reaches same phase in $t_{1}$ not in $t_{2}$, but in a closer point $t_{2}+\Delta t_{\rho}$ where $\Delta t_{\rho}$ is positive or negative, and $\Delta t_{\rho}=0$ if (3.9) hold.
-Such $\Delta t_{\rho}$ become smaller in absolute value, and, also the amplitude of oscillation evidenced in (3.3), decreases as $p_{max}$ ( and, cosequently $\rho$ ) grows.
-So the error terms at $\Delta t$ extrema must go to zero as $p_{max}\rightarrow\infty$, even if extremely slowly, because all $\ln(\rho)\times t$ are different.
+So the error terms, $\Delta R$, at $\Delta t$ extrema become very small as $p_{max}\rightarrow\infty$, even if extremely slowly, because for different $\rho_{k}$, big enough, we have to sum vectors with different amplitudes but phase almost uniformily distributed in $0,2\pi)$ interval for $t_{1..2}\ln(\rho)$, at least, but here ( differenly then in [15] ), also for the equidistribution of characters , see (3.19 ) .
 In order to improve convergence , pointing to bigger addends of (3.4), (i.e. at $\rho$ close to $p_{max}$ ), we can choose:
@@
 $L(s,\chi,p_{max})+R(s,\chi,p_{max})$ for $n$ and $\rho$ close to $p_{max}$ because phase difference in $\Delta t$ is: $\ln(\rho)\Delta t=\ln(p_{max})\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)\ \frac{2\pi}{\ln(p_{max})}=2\pi\left(1+\frac{\ln(r)}{\ln(p_{max})}\right)$ , and for $p_{max}\rightarrow\infty$ it tends to $2\pi$.
-So as concern these small $r$ terms we have (considering mean values in $\Delta t$ interval):
+So as concern these, same $\rho$, small $r$ terms we have (considering mean values in $\Delta t$ interval):
 (3.11) $\left.\Delta\angle[EP(s.\chi,p_{max})]\right|_{\Delta t}\rightarrow\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}\ \ ;$
 where $p^{*}$ is whatever big.
+The huge amount of terms that are not included in this cancellation (i.e .
+$\frac{\ln(r)}{\ln(p_{max})}$ not so small , and with $\rho=p_{max}\times r$) have a smaller amplitude and with phases distributed in $0,2\pi)$ interval.
 So condition (3.10) is an hint to convergence at lower $p_{max}$ while the argument stemming from identity (3.7) can be applied to generic $p_{max}>>p^{*}$ condition, but, it requires higher $p_{max}$ values to reach a comparable convergence.
 In a simplistic view we could say:
-(3.3) is mainly an oscillation with period close to $\frac{2\pi}{\ln(p_{max})}$, if we take $\Delta t=m\times\frac{2\pi}{\ln(p_{max})}\ ;\ m\in N$ we filter away the main part of (3.3) .
+(3.4) is mainly an oscillation with period close to $\frac{2\pi}{\ln(p_{max})}$, if we take $\Delta t=m\times\frac{2\pi}{\ln(p_{max})}\ ;\ m\in N$ we filter away the main part of (3.4) for each $\rho=p_{max}\times r$, when $\frac{\ln(r)}{\ln(p_{max})}$ is small.
+So we have cancellation of main part of $R$.
 Notice that the mean phase variation in $\Delta t$ is :
 (3.12) $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}$
+Conclusions : we must be cautious to affirm that $\L (s.\chi_{primitive})=EP(t,\epsilon,p_{max}\rightarrow\infty)$ in critical strip.
+Surely where $\L (s.\chi_{primitive})=0$ it is not true as here $R(t,\epsilon,\chi,p_{max})$, i.e.
+(3.4) is clearly not zero.The sum $R(t,\epsilon,\chi,p_{max})$ contains all the terms of $L(s,\chi)$ from $p_{max}<\rho\leq\rho_{max}$ but the integer containing the primes greater than $p_{max}$ and their powers.
+If we chose a whatsoever succession of $p^{k}_{max}$, like in (3.3 ) we have that the successive $R(t,\epsilon,\chi,p_{max}^{k})$ must be decorrelated as a noise-like succession, so that a built up is not possible for $p\rightarrow\infty$.
+It is known that $\sum_{p_{1}}^{\infty}\frac{1}{p^{\gamma}}\rightarrow\infty\ ,\forall p_{1}$, and $\gamma=1$ ( so much more for $\gamma<1$ ).
+But in (3.4), and, in (3.7), the term $\cos(\ln(\rho)t)-i\sin\cos(\ln(\rho)t)]\frac{\chi(\rho)}{\rho^{\Re[s]}}$ may impose a remarkable cancellation.
+On the other hand, we are interested in (3.12), i.e.
+not a simple difference, but, a phase difference.
+We will see that $|\Delta L|$ is always big meanwhile, for $\Delta t$ fixed, $|\Delta R|$, (3.7), is rather small due to cancellations in bigger $R$ terms by (3.10) choice, and by general cancellation mechanism in interval $\Delta t$, and, we conclude that $\angle[\Delta L+\Delta R]\approx\angle[\Delta L]$ when $\angle[\Delta L]>>0$ ( i.e.
+close to $s^{*}:L(s^{*},\chi)=0$) or $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}>0$ (i.e.
+on critical line far from zeros, see (A.14 ) , and, Lemma 2).
+So, where the above conditions are verified, we have some merit to consider :
+(3.13) $\frac{\left.\Delta\angle[L(s.\chi)]\right|_{\Delta t}}{\Delta t}=\frac{\left.\Delta\angle[EP(s.\chi)]\right|_{\Delta t}}{\Delta t}\ \ \ \ ;\ \ \Delta t=\frac{2\pi}{\ln(p^{*})}\ ;\ \ p_{max}=(p^{*})^{m};\ m\in N\ ,\ m\rightarrow\infty$
+Further examples will confirm this position.
 3.2. Application of Euler Product in critical strip
@@
 We neglect, for the moment, double zeros.
-To be specific $\xi(s,\chi)$ is equal ( [6, p.
+To be specific $\xi(s,\chi)$ is equal ( [7, p.
 82]) to the infinite Hadamard product, that has single factors like
@@
 Notice this discontinuity is not present in (2 ), where total phase is not used.
-Now let us see how to evaluate phase variations of expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$ using (3.9 ).
+Now let us see how to evaluate phase variations of expression $\angle[L(s.\chi_{primitive})]=\angle\left[\sum_{n=1}^{\infty}\ \chi(n)/n^{-1}\right]$ using (3.11 ).
 In order to trade resolution along $t$ with computational burden it is convenient to address to mean phase variation in an interval $(t_{1},t_{2})$ .
@@
 Notice that for $p$ big ( i.e.
-$p^{1/2+\epsilon}>>1\geq\left|\cos(\ln(p)t-\angle[\chi(p)])\right|\ \ \$), the content of braces summed in (3.13 tends to be:
+$p^{1/2+\epsilon}>>1\geq\left|\cos(\ln(p)t-\angle[\chi(p)])\right|\ \ \$), the content of braces summed in (3.16 tends to be:
 (3.17) $\left\{\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right\}_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\rightarrow$
@@
 […]
-We want to point out at $h$-prime classes in (3.13). If we limit to a finite $p_{max}$ we can write (3.13 ) as the sum of the $\phi(q)$ sums involving the $h$-prime classes:
+We want to point out at $h$-prime classes in (3.16). If we limit to a finite $p_{max}$ we can write (3.16 ) as the sum of the $\phi(q)$ sums involving the $h$-prime classes:
 (3.21) $\frac{\ln(p^{*})}{2\pi}\sum_{h<q\ :\ gcd(h,q)=1}\left\{\sum_{p<p_{max}\ :\ p\equiv h(mod\ q)}\left[\arctan\left(\frac{\sin(\ln(p)t-\angle[\chi(p)])}{p^{1/2+\epsilon}-\cos(\ln(p)t-\angle[\chi(p)])}\right)\right]_{t-\frac{\pi}{\ln(p^{*})}}^{t+\frac{\pi}{\ln(p^{*})}}\right\}$
@@
 (3.25) $\ln\left[\sqrt{\frac{t\times q}{2\pi}}\right]+\left\{\left[\frac{\Delta\angle[L_{EP}(s,\chi_{primitive})]}{\Delta t}\right]_{p_{max}\rightarrow\infty}\right\}_{p^{*}\rightarrow\infty}\rightarrow\left[\frac{\partial\angle[\xi(t,\epsilon,\chi)]}{\partial t}\right]_{\epsilon=0}=0$
-PROOF. From Lemma 2, appendix A and (3.13 ) follows (3.22 ). End of proof.
+PROOF. From Lemma 2, appendix A and (3.16 ) follows (3.25 ). End of proof.
 In half-plane $\Re[s]>0$ (see also figg. 5, 4, and , 3)) , we can remark that:
@@
 See (2), where, only phase difference between vector of real and imaginary components with respect to vector of their $t$-derivative matters.
-  • (3) $\frac{\Delta\angle[L_{EP}(t,\epsilon,\chi)]}{\Delta t}$ instead is defined everywhere but points $(\epsilon\leq\epsilon^{*}_{k}\ ,t=t^{*}_{k})$ , see (3.13 ). Here is used total phase.
+  • (3) $\frac{\Delta\angle[L_{EP}(t,\epsilon,\chi)]}{\Delta t}$ instead is defined everywhere but points $(\epsilon\leq\epsilon^{*}_{k}\ ,t=t^{*}_{k})$ , see (3.16 ). Here is used total phase.
 3.5. LEMMA 6: Sign estimate of $\frac{\partial^{2}\angle[L(s,\chi_{primitive})]}{\partial\epsilon\partial t}$
@@
 Then,for (3.25) we have that (3.16) evaluated in $\epsilon=0$ is tending to $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]<0$ as $(p_{max}\rightarrow\infty)p^{*}\rightarrow\infty]$’, in order to verify Lemma 2.
-We affirm that the same (3.13), evaluated for small $\epsilon>0$ is always negative but with lower absolute value, so that (3.22 ) has a positive increment for small $\epsilon>0$.
+We affirm that the same (3.16), evaluated for small $\epsilon>0$ is always negative but with lower absolute value, so that (3.25 ) has a positive increment for small $\epsilon>0$.
 i,e (at least for odd primitive characters, see (3.24) ):
@@
 (3.31) $O^{+}_{k,h}(\epsilon)_{\int dLi}=\left|\frac{\ln(p^{*})}{\pi\phi(q)}\int_{x>x_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{x<x^{\prime}_{0t}(k,h)}\frac{\cos[\ln(x)t-\angle[\chi(h)]]\sin[\pi\ln(x)/\ln(p^{*})]}{\sqrt{x}x^{\epsilon}}dLi(x)\right|$
-Negative contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}$):
+Negative contribution of (3.16) (and of final result) is given by the $h$-class primes (weighted by (3.29 ) function) inside the analogous intervals $\Delta y_{k,h}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}$):
 (3.32) $O^{-}_{k,h}(\epsilon)_{\sum}=\left|\frac{\ln(p^{*})}{2\pi}\sum_{p>y_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{p<y^{\prime}_{0t}(k,h)}F(p,\chi(p),\epsilon)\right|$
@@
 (3.34) $O^{-}_{k,h}(\epsilon)_{\int dLi}=\left|\frac{\ln(p^{*})}{\pi\phi(q)}\int_{x>x^{\prime}_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{x<x_{0t}(k+1,h)}\frac{\cos[\ln(x)t-\angle[\chi(p)]]\sin[\pi\ln(x)/\ln(p^{*})]}{\sqrt{x}x^{\epsilon}}dLi(x)\right|$
-Positive contribution of (3.13) (and of final result) is given by the $h$-class primes (weighted by (3.26 ) function) inside the analogous intervals $\Delta y_{k,h}^{\prime}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}^{\prime}$)
+Positive contribution of (3.16) (and of final result) is given by the $h$-class primes (weighted by (3.29 ) function) inside the analogous intervals $\Delta y_{k,h}^{\prime}$ ( that for $k$ big tends to the previous one $\Delta x_{k,h}^{\prime}$)
 (3.35) $O^{+}_{k,h}(\epsilon)_{\sum}=\left|\frac{\ln(p^{*})}{2\pi}\sum_{p>y^{\prime}_{0t}(k,h)\ :\ p\equiv h(mod\ q)}^{p<y_{0t}(k+1,h)}F(p,\chi(p),\epsilon)\right|$
@@
 In other words, the sign contributions of final result, i.e.
-(3.13 ) minus (3.19), or , (3.20), in intervals $\Delta x,\Delta x^{\prime},\Delta y,\Delta y^{\prime}$, is as that stated in ( 3.29), (3.32) , and, (3.28), (3.31).
+(3.16 ) minus (3.22), or , (3.23), in intervals $\Delta x,\Delta x^{\prime},\Delta y,\Delta y^{\prime}$, is as that stated in ( 3.32), (3.35) , and, (3.31), (3.34).
 Instead where $m$, in (3.36), is odd, then, in every interval above the sign changes concurrently.
@@
 (3.39) $\sum_{+}^{p_{max}}(\epsilon)=\sum_{k}\sum_{h}O^{+}_{k,h}(\epsilon)_{\sum}+O^{+}_{k,h}(\epsilon)_{\int dLi}$
-We stress again, (3.34 ) is computed without changing the order given by $p$ or $x=y$, like in (3.18 ) , i.e.
+We stress again, (3.37 ) is computed without changing the order given by $p$ or $x=y$, like in (3.21 ) , i.e.
 partial sums are unaffected.
-Limit difference in (3.34 ) is in form: $\infty$-$\infty$, and, we know that if $|t-t_{k}^{*}|>\frac{2\pi}{\ln(p^{*})}$ limit result is finite .
-In other words: positive quantities: $O^{+}_{kh}(\epsilon)_{\int dLi}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\int dLi}$,and , $O^{+}_{k,h}(\epsilon)_{\sum}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\sum}$ like in ( 3.29), and (3.32) , and, (3.28), (3.31) are only “place-holders” filled according to progress of $x=y<p_{max}$, and with chosen sign following the $m$ parity in (3.33).
+Limit difference in (3.37 ) is in form: $\infty$-$\infty$, and, we know that if $|t-t_{k}^{*}|>\frac{2\pi}{\ln(p^{*})}$ limit result is finite .
+In other words: positive quantities: $O^{+}_{kh}(\epsilon)_{\int dLi}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\int dLi}$,and , $O^{+}_{k,h}(\epsilon)_{\sum}\ \ ,\ \ O^{-}_{k,h}(\epsilon)_{\sum}$ like in ( 3.32), and (3.35) , and, (3.31), (3.34) are only “place-holders” filled according to progress of $x=y<p_{max}$, and with chosen sign following the $m$ parity in (3.36).
 Consider now the ratios :
 (3.40) $\frac{O^{-}_{k,h}(\epsilon^{\prime})_{\sum}}{O^{+}_{k,h}(\epsilon^{\prime})_{\int dLi}}\ \ \ \ \ \ ;\ \ \ \ \ \ \frac{O^{+}_{k,h}(\epsilon^{\prime})_{\sum}}{O^{-}_{k,h}(\epsilon^{\prime})_{\int dLi}}$
-for each $h$ and for $k$ big, defined respectively in intervals $\Delta y_{k,h}\rightarrow\Delta x_{k,h}$ ( see (3.28) and (3.29 ) ) , and in intervals $\Delta y_{k,h}^{\prime}\rightarrow\Delta x_{k,h}^{\prime}$ ( see (3.31) and (3.32) ) .
-LEMMA 7: $\forall h,k$ as $k\rightarrow\infty$ both (3.37 ) tend to unity.
-PROOF. PNT ratio in small intervals [8, p. 23] assures that:
+for each $h$ and for $k$ big, defined respectively in intervals $\Delta y_{k,h}\rightarrow\Delta x_{k,h}$ ( see (3.31) and (3.32 ) ) , and in intervals $\Delta y_{k,h}^{\prime}\rightarrow\Delta x_{k,h}^{\prime}$ ( see (3.34) and (3.35) ) .
+LEMMA 7: $\forall h,k$ as $k\rightarrow\infty$ both (3.40 ) tend to unity.
+PROOF. PNT ratio in small intervals [9, p. 23] assures that:
 (3.41) $\pi(x+\Phi(x))-\pi(x)\sim\frac{\Phi(x)}{\ln(x)}\ \ with\ \ x\rightarrow\infty\ \ ,if\ \Phi(x)\geq x^{7/12-\epsilon(x)}$
@@
 for $x$ big enough the distribution of primes in equal-sign intevals $\Delta x_{k,h}\rightarrow\Delta y_{k,h}$, and, for (3.19) also $h$-class primes, can be close to theorethical one as we want.
-In [17, p.
+In [18, p.
 65] it is dealt about the sum of prime sampled functions: “in the neighborhood of the number $x$ the average density of the primes is $1/\ln(x)$.
@@
 Besides when $p_{max}$ satisfy (3.36) with odd $m$ then in ratio (3.45) we add $O^{-}_{k,h}(\epsilon^{\prime\prime})_{\sum}+O^{-}_{k,h}(\epsilon^{\prime\prime})_{\int dLi}$ at numerator and $O^{+}_{k,h}(\epsilon^{\prime})_{\int dLi}+O^{+}_{k,h}(\epsilon^{\prime})_{\sum}$ at denominator, and, we have always the ratio beween all positive versus all negative contributions.
-Now we want the ratio of all positive contributions to final result (3.34 ) at different $\epsilon$ value.
+Now we want the ratio of all positive contributions to final result (3.37 ) at different $\epsilon$ value.
 In detail : $\forall h<q\ ,gcd(h,q)=1$, $\epsilon^{\prime\prime}>0$, and $\epsilon^{\prime}=0$:
@@
 (3.47) $\frac{\sum_{-}^{p_{max}}(\epsilon)}{\sum_{-}^{p_{max}}(0)}=\frac{\sum_{k}\sum_{h}O^{-}_{k,h}(\epsilon)_{\sum}+O^{-}_{k,h}(\epsilon)_{\int dLi}}{\sum_{k}\sum_{h}O^{-}_{k,h}(0)_{\sum}+O^{-}_{k,h}(0)_{\int dLi}}=\rho^{-}(\epsilon,p_{max})\ \ \ <\ \ 1;\ \ \epsilon>0$
-(3.44 ) and (3.43 ), if $\epsilon>0$ is fixed and $p_{max}\rightarrow\infty$ goes to zero, because if $\epsilon>0$ and $k$ i.e.
+(3.47 ) and (3.46 ), if $\epsilon>0$ is fixed and $p_{max}\rightarrow\infty$ goes to zero, because if $\epsilon>0$ and $k$ i.e.
 $p_{max}\rightarrow\infty$, then $p^{-\epsilon}\rightarrow 0$.
@@
 (3.48) $\sum_{+}^{p_{max}<\infty}(\epsilon>0)<\sum_{+}^{p_{max}<\infty}(\epsilon=0)\ \ \ \ \ ;\ \ \sum_{-}^{p_{max}<\infty}(\epsilon>0)<\sum_{-}^{p_{max}<\infty}(\epsilon=0)$
-Besides the indeterminate limit (3.34 ), of $\infty$-$\infty$ form, for $\epsilon>0$ ($p^{*}$ is big and fixed), gives :
+Besides the indeterminate limit (3.37 ), of $\infty$-$\infty$ form, for $\epsilon>0$ ($p^{*}$ is big and fixed), gives :
 (3.49) $\sum_{-}^{p_{max}\rightarrow\infty}(\epsilon)=\sum_{+}^{p_{max}\rightarrow\infty}(\epsilon)+X(\epsilon,p_{max})\ \ ;\ \ X(\epsilon=0,p_{max}\rightarrow\infty)=-\left[\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi_{primitive})]}{\Delta t}\right]_{p_{max}\rightarrow\infty}$
@@
 As $p_{max}$ grows, $\rho^{+}(\epsilon,p_{max})\rightarrow\rho^{-}(\epsilon,p_{max})$ for suitable small $\epsilon>0$, both approach 1, but, from below.
-See (3.47) and ( 3.48).
+See (3.50) and ( 3.51).
 So $X(\epsilon>0,p_{max})=\rho^{\pm}(\epsilon,p_{max})X(\epsilon=0,p_{max})$ for $p_{max}$ big enough.
@@
 So we can state :
-Corollary of Lemma 6. From (3.46 ) and from (3.13) we can write for $\epsilon\rightarrow 0^{+}$:
+Corollary of Lemma 6. From (3.49 ) and from (3.16) we can write for $\epsilon\rightarrow 0^{+}$:
 (3.54) $\frac{1}{\epsilon}\left[\frac{\Delta\angle[L_{EP}(t,\epsilon>0,\chi)]}{\Delta t}-\frac{\Delta\angle[L_{EP}(t,\epsilon=0,\chi)]}{\Delta t}\right]_{\ p_{max}\rightarrow\infty,\ p^{*}\ big}\rightarrow\left[\frac{\partial^{2}\angle[L(s,\chi_{primitive}))]}{\partial\epsilon\partial t}\right]_{\epsilon=0}\geq 0$
@@
 $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$ for $\epsilon=0$), so, for $p^{*}$ big the difference of the mean in the intervals (3.6) is close to the function on the right in (3.54).
-Besides (3.23) holds for every big $p^{*}$, so, we can work on (3.46 ).
+Besides (3.26) holds for every big $p^{*}$, so, we can work on (3.49 ).
 For [2, p.
@@
 $\rho^{\pm}(\epsilon^{\prime},p^{\prime}_{max})>1$.
-But this is not possible for $p^{\prime}_{max}$ big, and $\forall p^{*}$, see (3.45 ), (3.47).
+But this is not possible for $p^{\prime}_{max}$ big, and $\forall p^{*}$, see (3.48 ), (3.50).
 See Lemma 6 (3.26).
@@
 $\mathcal{L}[\eta(s,\chi)]=\mathcal{L}[\xi(s,\chi)]=|\xi(s,\chi))|^{2}\times\frac{\partial\angle[\xi(s,\chi))]}{\partial t}=|\eta(s,\chi))|^{2}\times\frac{\partial\angle[\eta(s,\chi))]}{\partial t}\ \ :\ \ \chi_{primitive}$
-From (3.22 ) , as for Lemma 2:
+From (3.25 ) , as for Lemma 2:
 $\left[\frac{\partial\angle[L(s,\chi_{primitive})]}{\partial t}\right]_{\epsilon=0}=-\ln\left(\sqrt{\frac{tq}{2\pi}}\right)\ ;\ \forall t\neq t^{*}_{k}$
@@
 So $|\Im[\xi(\frac{1}{2}+\epsilon+it)]|$ grows with $\epsilon$ far from $\Re[\xi(\frac{1}{2}+it)]$ extremal points.
-We do not say that the set of points of $(t,\epsilon)$ plane where $\Im[\xi(s)]=0$ are perpendicular to critical line at $t_{0}$ where $\Re[\xi(\frac{1}{2}+it_{0})]$ is extremal, but, the truth is not far from it [12, p.
-3] , and [13, p.
+We do not say that the set of points of $(t,\epsilon)$ plane where $\Im[\xi(s)]=0$ are perpendicular to critical line at $t_{0}$ where $\Re[\xi(\frac{1}{2}+it_{0})]$ is extremal, but, the truth is not far from it [13, p.
+3] , and [14, p.
 9,fig 6].
@@
 $\frac{\partial\angle\left[\left(\frac{\pi}{q}\right)^{-\frac{s+\alpha}{2}}\Gamma\left(\frac{s+\alpha}{2}\right)\right]}{\partial t}=\ln\left(\sqrt{\frac{tq}{2\pi}}\right)+O(t^{-2})$, and $O(t^{-2})\rightarrow+(-1)^{\alpha+1}\frac{\epsilon}{4t^{2}}$ for $\epsilon\approx 0$ ($\alpha=0$ even, and $\alpha=1$ odd characters).
-Here we find same results of appendix A without Stirling formula like in expression (A.5). The purpose is to evaluate (A.13) in the $t=0$ neighborhood. The starting point is the formula ([7, p. 8]) )
+Here we find same results of appendix A without Stirling formula like in expression (A.5). The purpose is to evaluate (A.13) in the $t=0$ neighborhood. The starting point is the formula ([8, p. 8]) )
 (4.1) $\Gamma(z)=\lim_{N\to\infty}\frac{N!\ \ (N+1)^{z}}{z(z+1)...(z+N)}$
@@
 (4.3) $\sum_{1}^{N\rightarrow\infty}\left\{\frac{1}{2n}-\left(\frac{1}{2n+1/2+\epsilon+\alpha}\right)\left(\frac{1}{1+\left(\frac{t}{2n+1/2+\epsilon+\alpha}\right)^{2}}\right)\right\}$
-Without Stirling formula as in (A.14 ) , the equivalent relation of $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]$ in (3.22 ) or (3.21), is (4.4).
+Without Stirling formula as in (A.14 ) , the equivalent relation of $\ln\left[\sqrt{\frac{tq}{2\pi}}\right]$ in (3.25 ) or (3.24), is (4.4).
 See fig.
@@
 can be used also locally if the zeros happen at $t$ values in which same situation (4.5) is locally met.
-The zeros of $L(s,\chi_{primitive})$ distribute like the peaks of the expression (3.13 ) that we know occurs exactly at the zero of $L(s,\chi_{primitive})$.
+The zeros of $L(s,\chi_{primitive})$ distribute like the peaks of the expression (3.16 ) that we know occurs exactly at the zero of $L(s,\chi_{primitive})$.
 Irrespective if RH is verified or not.
@@
 256].
-Table loaded from [11] and adapted for (3.13 ) .
+Table loaded from [12] and adapted for (3.16 ) .
 q Odd chrs Suff. Cond. (4.5) closest odd chrs correlation peak distance from $t=0$
@@
 Figure 3.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in not principal character case $q=5$ ( see tab.
-1) , with $p_{max}=p^{*}=158\times 10^{6}$ ( here are used the first 8868881 primes in (3.13 ) ).In wolfram demostration project - Dirichlet L-Functions and their zeros we can compute, by zeroing real and imaginary parts, the zeros closer to $t=0$ of what here are called X1 and X3, the result is: $\pm 9.443$, $\pm 8.457$, $\pm 6.184,\ \ \pm 4.133\$.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in not principal character case $q=5$ ( see tab.
+1) , with $p_{max}=p^{*}=158\times 10^{6}$ ( here are used the first 8868881 primes in (3.16 ) ).In wolfram demostration project - Dirichlet L-Functions and their zeros we can compute, by zeroing real and imaginary parts, the zeros closer to $t=0$ of what here are called X1 and X3, the result is: $\pm 9.443$, $\pm 8.457$, $\pm 6.184,\ \ \pm 4.133\$.
 Very comparable with results in figure.
 Figure 4.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=5$, with $p_{max}=(p^{*})^{3}=45.8\times 10^{6}$ ( here are used the first 2763823 primes in (3.13 ) ).The peaks are lower and fatter with respect to fig.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=5$, with $p_{max}=(p^{*})^{3}=45.8\times 10^{6}$ ( here are used the first 2763823 primes in (3.16 ) ).The peaks are lower and fatter with respect to fig.
 3, but, the inter-peaks behavior is much more close to $\ln\left(\sqrt{\frac{qt}{2\pi}}\right)$, or, better, to (4.4), that follows different behavior close to $t=0$ if $\alpha=0$ (even chrs) or $\alpha=1$ (odd chrs).
 Figure 5.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=3$ and $q=4$, with $p_{max}=(p^{*})^{3}=158\times 10^{6}$ ( here are used the first 8868881 primes).For $q=3$, and, $\chi_{1}(n)=(0,1,-1)$ or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=3$ and $q=4$, with $p_{max}=(p^{*})^{3}=158\times 10^{6}$ ( here are used the first 8868881 primes).For $q=3$, and, $\chi_{1}(n)=(0,1,-1)$ or $\angle[\chi_{1}(n)]=(undefined,0,\pi)$ to comply with format of tab.
 1.
@@
 4.
-In this case compare with fig.
-14.
+In this case compare the inter-peaks level with fig.
+15 and asymptotic formula (A.14 ).
 Putting all pieces together we can state that $\xi(s,\chi_{odd\ primitive})$, is RH compliant $\forall t$ and $\forall q$ .
@@
 $\chi_{5}(n)$ 0 11 $w^{2}$ 0 w-w w-w 0 $w^{2}$ 1-1
-Table 4. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=9$, $w=e^{i\pi/3}$. I Table loaded from [11] .
+Table 4. Characters components for $L(s,\chi(n))$ .Arithmetic congruence modulus $q=9$, $w=e^{i\pi/3}$. The table is loaded from [12] .
 Figure 6.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab.
 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes ).
 Figure 7.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ (see tab.
 3), and , with $p_{max}=(p^{*})=5.8\times 10^{6}$ ( here are used the first 400038 primes).
 Figure 8.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) flipped vertically.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) flipped vertically.
 So Lemma 2 is verified.
 Figure 9.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).
+The expression (3.16 ) in character case $q=13$, and, for $\chi_{10}$ has been computed with $\epsilon=0$ and $\epsilon=0.1$.
+Notice that with greater $\epsilon$ the curve is close to $t$ axis as predicted in Lemma 6 .
+See (3.26 ).
+The choices are: $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).
+Figure 10.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=13$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).
 Notice that inter-peaks level is the same of fig, (15) flipped vertically.
@@
 Figure 11.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically.
 So Lemma 2 is verified.
 Figure 12.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (2) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(3/\pi)}{2}$ and afterward flipped vertically.
 So Lemma 2 is verified.
 Figure 13.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=p^{*}=5.8\times 10^{6}$ ( here are used the first 400038 primes).Notice that inter-peaks level tend to be the same of fig, (15) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically.
 So Lemma 2 is verified.
 Figure 14.
-The peaks symmetries of (3.13 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (14) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically.
+The peaks symmetries of (3.16 ) with $\epsilon=0$, respect to real axis, in character case $q=9$ and , with $p_{max}=(p^{*})^{3}=58\times 10^{6}$ ( here are used the first 3450491 primes).Notice that inter-peaks level is the same of fig, (15) with a shift by the quantity $\frac{\ln(9/\pi)-\ln(5/\pi)}{2}$ and afterward flipped vertically.
 So Lemma 2 is verified.
@@
 If we consider the factor (s1)(s-1) attached to $\zeta(s)$ like in (A.1), we can use (A.4) again putting $\alpha_{1}=0\ ,\ \alpha=2\ ,\ q=1$.
-As $z+1=\frac{s+\alpha}{2}\rightarrow z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$, Stirling formula ( [7, p.
+As $z+1=\frac{s+\alpha}{2}\rightarrow z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$, Stirling formula ( [8, p.
 109,112]) allows us to write (A.4) for each of the 3 choices of $(\alpha,\alpha_{1})$:
@@
 And $\Re_{2}$ is the real part.And finally let us call $\Re_{3}+i\Im_{3}=\left(\sum_{k=1}^{K-1}\frac{B_{2k}}{2k(2k-1)z^{2k-1}}\right)+R_{2K}(z)$.
-The $B_{j}$ are the Bernoulli numbers ( [7, p. 11] ), that vanish for odd $j$ while:
+The $B_{j}$ are the Bernoulli numbers ( [8, p. 11] ), that vanish for odd $j$ while:
 $B_{2}=\frac{1}{6}\quad;\quad B_{4}=-\frac{1}{30}\quad;\quad B_{6}=\frac{1}{42}\quad;\quad B_{8}=-\frac{1}{30}\quad;\quad B_{10}=\frac{5}{66}\quad;\quad B_{12}=-\frac{691}{2730}\quad$...
@@
 (A.6) $|R_{2K}(z)|<\left(\frac{B_{2K}}{2K(2k-1)z^{2K-1}}\right)\frac{1}{\left[\cos\left(\frac{\arg(z)}{2}\right)\right]^{2K}}$
-where $\arg(z)$ is taken in the interval: $\pi<\arg(z)<\pi$ (see [7, p. 112].
+where $\arg(z)$ is taken in the interval: $\pi<\arg(z)<\pi$ (see [8, p. 112].
 In (A.5) we may put : $z=\frac{2\epsilon+2\alpha-3}{4}+i\ \frac{t}{2}$ , so, leaving apart only $\Im_{3}$ for now, we have:
@@
 (B.4) $\frac{\ln(p^{*})}{4\pi}\sum_{p<p_{max}\ :\ gcd(p,q)=1}\left(\frac{\sin(2(\ln(p)t_{2}-\angle[\chi(p)]))-\sin(2(\ln(p)t_{1}-\angle[\chi(p)]))}{[p^{1/2+\epsilon}-\cos(\ln(p)t_{2}-\angle[\chi(p)])][p^{1/2+\epsilon}-\cos(\ln(p)t_{1}-\angle[\chi(p)])]}\right)$
-It is apparent that the difference between computation (3.13 ), upon substitution of (3.14) with (3.15) is absolutely convergent for sure from $\epsilon>0$ and conditionally convergent for $\epsilon=0$.
-Proof. Let us look at (3.27), and, consider only one $h$-class primes till $p_{max}$ ( as in (3.18 ).
+It is apparent that the difference between computation (3.16 ), upon substitution of (3.17) with (3.18) is absolutely convergent for sure from $\epsilon>0$ and conditionally convergent for $\epsilon=0$.
+Proof. Let us look at (3.30), and, consider only one $h$-class primes till $p_{max}$ ( as in (3.21 ).
 Each equal sign interval (see (3.30)) now $\Delta p=\frac{\pi}{2t}p$ alternating sign, in (B.4) is divided by $\approx p$.
@@
 This seems confirmed by figg.
-17, 18, and, 19.
+18, 19, and, 20.
 Figure 17.
@@
 Equation (1.1) does not hold for $0<\Re(s)<1$, but, subsection 3.1 shows a way in order to exploit Euler product for variations, in particular phase variations, on small $t$-intervals.
-On the other hand ( even if it is not an argument) it seems to me worth to draw the attention on figg, 3 through 13 where it is apparent that (3.9), at least numerically, works rather well even with only $m=3$.
+On the other hand ( even if it is not an argument) it seems to me worth to draw the attention on figg, 3 through 14 where it is apparent that (3.11), at least numerically, works rather well even with only $m=3$.
 In conclusion it seems to me that subsection 3.1 is a valid base for the developments of section 3.
@@
     ANSW.
-I agree with the statement, but it is not applicable.The “manipulation” of terms in (3.34 ) (or in others occasions (3.1), (3.18) and (3.20) ) is applied only at addends with index less than the running index.
+I agree with the statement, but it is not applicable.The “manipulation” of terms in (3.37 ) (or in others occasions (3.1), (3.21) and (3.23) ) is applied only at addends with index less than the running index.
 So on a set of finite numbers.
@@
     If we take the sum till $J_{max}$ and then we partition the addends with $j\leq J_{max}$ in positive $\sum^{+}_{J_{max}}$ and negatives $\sum^{-}_{J_{max}}$, then we can write $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$, and, we do not alter final result $\ln(2)$ for $J_{max}\rightarrow\infty$ transforming the limit in something like $\infty$-$\infty$.
-It is exactly the “manipulation” in expression : (3.34 ).
+It is exactly the “manipulation” in expression : (3.37 ).
 While in (3.1), in (3.21), and, in (3.23), we manipulate to highlight some structure of the series, but , with indexes less than the running index .
@@
 So observation does not apply.
-  • The argument in equations ( 3.43 ) through (3.50) rests on the idea that because each term in the series for $\epsilon>0$ is smaller in magnitude than the corresponding term for $\epsilon=0$ (due to the factor of $p^{-\epsilon}$), the overall sum must also be smaller in magnitude.
+  • The argument in equations ( 3.46 ) through (3.53) rests on the idea that because each term in the series for $\epsilon>0$ is smaller in magnitude than the corresponding term for $\epsilon=0$ (due to the factor of $p^{-\epsilon}$), the overall sum must also be smaller in magnitude.
 This reasoning is incorrect.
     ANSW.
-( 3.43 ) is a ratio not of the whole sum, but only of the positive contributions at different $\epsilon$ values.
-Each positive contribution decreases with increasing $\epsilon$, the same holds for their sum, so this ratio ( 3.43 ) also $\forall p_{max}<\infty$ is less than 1 and also (3.45 ) are completely justyfied.
+( 3.46 ) is a ratio not of the whole sum, but only of the positive contributions at different $\epsilon$ values.
+Each positive contribution decreases with increasing $\epsilon$, the same holds for their sum, so this ratio ( 3.46 ) also $\forall p_{max}<\infty$ is less than 1 and also (3.48 ) are completely justyfied.
 The same holds for negative contributions.
@@
 The computation takes in consideration correct contributions.
-  • Analyzing the sum of positive terms in isolation from the sum of negative (like in (3.45 ) terms is not a valid operation
-    ANSW Generally speaking I can agree but (see example two bullets above) $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$ tends exacly to the limit.
+  • Analyzing the sum of positive terms in isolation from the sum of negative (like in (3.48 ) terms is not a valid operation
+    ANSW Generally speaking I can agree but (see example two bullets above) $S_{J_{max}}=\sum^{+}_{J_{max}}-\sum^{-}_{J_{max}}$ tends exacly to the limit, provided the operation is carried on a finite terms set.
 So I do not understand the sentence: it“is not a valid operation”.
-  • algebraic manipulations from ( 3.43 ) through (3.50) effectively treat the series as if it were a sum of positive terms, ignoring the crucial role of cancellation.
+  • algebraic manipulations from ( 3.46 ) through (3.53) effectively treat the series as if it were a sum of positive terms, ignoring the crucial role of cancellation.
 This is a classic error when dealing with conditional convergence
@@
 A valid argument would require the use of an explicit formula for the distribution of primes, complete with error terms, and demonstrate that these error terms do not affect the final result.
-    ANSW In Lemma 4 it is used the function $Li(x)$, but, with whatsoever well behaved function the result of (3.19 ) is null because it depends from (3.17).
+    ANSW In Lemma 4 it is used the function $Li(x)$, but, with whatsoever well behaved function the result of (3.22 ) is null because it depends from (3.20).
 So in Lemma 4 there is no reference to (PNTAP).
 The rationale of (3.23) is that of an algebraic trick to insert, with zero effective effect, the integral of PNT in (3.16).
-The difference in (3.20) means nothing because we are subtracting zero (i.e.(3.19 )) from (3.13).
+The difference in (3.23) means nothing because we are subtracting zero (i.e.(3.22 )) from (3.16).
 But it is useful in Lemma 6 to build not differences between primes distribution $\pi(x)$ and $Li(x)$ but ratios between functions involving $\pi(x)$ and $Li(x)$.
@@
 Let us say they are a necessary condition not a sufficient one.
-However I think that the matching of the $t_{peak}$ of the peaks of (3.13) with the zeros of $L(s,\chi)$, and the matching of the floor with the computed one by (A.14 ) and (4.4) (see on figg, 3 through 13) is an interesting fact.
+In other words, for example, it makes sense to say that fig.
+9 is compatible, i.e verifies, Lemma 6.
+Of course we cannot say it prove it.
+However I think that the matching of the $t_{peak}$ of the peaks of (3.16) with the zeros of $L(s,\chi)$, and the matching of the floor with the computed one by (A.14 ) and (4.4) (see on figg, 3 through 14) is an interesting fact.
   • The attempt to link mathematical problems to other fields like physics is not inherently frivolous, but, the lack of peer-reviewed validation makes it a speculative effort rather than a serious approach ANSW The reference to a kinematic quantity as angular momentum is only a short way to describe the method I explored.
@@
   • The term “spectrum of primes” is not a rigorously defined concept ANSW Perhaps you are right.
-I borrowed in ([14]) this term by [16].
+I borrowed this term from [17], and stressed it in ([15]) .
 But anyhow a formula that gather zeros of L-functions and primes ( as (3.16) ), seems interesting.